### v3

# PropMolFlow: Property-guided Molecule Generation with Geometry-Complete Flow Matching

Abstract
Molecule generation is advancing rapidly in chemical discovery and drug design. Flow matching methods have recently set the state of the art (SOTA) in unconditional molecule generation, surpassing score-based diffusion models. However, diffusion models still lead in property-guided generation. In this work, we introduce PropMolFlow, a novel approach for property-guided molecule generation based on geometry-complete SE(3)-equivariant flow matching. Integrating five different property embedding methods with a Gaussian expansion of scalar properties, PropMolFlow achieves competitive performance against previous SOTA diffusion models in conditional molecule generation across various properties while preserving the stability and validity of the generated molecules, consistent with its unconditional counterpart. Additionally, it enables faster sampling speed with significantly fewer time steps compared to baseline models. We highlight the importance of validating the properties of generated molecules through DFT calculations performed at the same level of theory as the training data. Specifically, our analysis identifies properties that require DFT validation and others where a pretrained SE(3) geometric vector perceptron regressors provide sufficiently accurate predictions on generated molecules. Furthermore, we introduce a new task to assess the model’s ability to propose molecules with underrepresented property values, assessing its capacity for out-of-distribution generalization. Our findings reveal shortcomings in existing structural metrics, which mistakenly validate open-shell molecules or molecules with invalid valence-charge configurations, underscoring the need for improved evaluation frameworks. Overall, this work paves the way for developing targeted property-guided generation methods, enhancing the design of molecular generative models for diverse applications.
Keywords Molecule Generation  $\cdot$ Property Guidance  $\cdot$ Flow Matching  $\cdot$ SE(3) Equivariant Graph Neural Networks
Introduction
Deep generative models show great promise to accelerate chemical discovery and drug design by direct statistical sampling of stable structures compared to resource-intensive physics-based approaches [1, 2]. The past few years have witnessed rapid development of generative methods for molecule generation, such as variational autoencoders (VAEs) [3, 4], generative adversarial networks (GANs) [5], autoregressive models [6], score-based diffusion models [7, 8], flow-based models [9, 10, 11] and language models [12, 13]. The predominant methods for molecule generation in 3D apply diffusion models combined with equivariant graph neural networks to generate molecules directly in 3D, providing a more accurate representation of molecular configurations compared to 1D SMILES and 2D graph-based representations [7, 14, 11, 8]. Flow matching, in particular, has emerged as a flexible alternative to diffusion models, providing greater control over the choice of prior distributions and probability paths, as well as enhanced sampling efficiency [15, 16, 17]. Its advantages have led to improved performance across various applications, including materials generation [18], protein design [19, 20], and molecule generation [21].
Seminal examples of diffusion models and flow matching for molecule generation treated molecules as point clouds and adopted a continuous representation of the atom types, leveraging E(3)-equivariant graph neural networks (EGNN) to learn the score function in diffusion models and the conditional velocity field in flow matching [7, 21]. However, the continuous representation of discrete molecular features, such as atom types and bond orders, fails to preserve their inherent discreteness, thus compromising their fidelity. Additionally, the EGNN architecture fails to capture molecular chirality, as an SE(3) generative process is required to account for chiral centers [22]. Recent studies also highlight the importance of incorporating bond orders in molecular graphs to improve the stability and validity of generated molecules [23, 24]. Building on these insights, recent diffusion models and flow matching approaches adopt a geometry-complete SE(3) generative process that incorporates bond orders into molecular graph representations, and utilize discrete diffusion or discrete flow matching for discrete molecular features such as atom types and bond orders [25, 26]. Notably, FlowMol integrates these advances to achieve SOTA performance in unconditional molecule generation [26]. However, its formulation remains limited to the unconditional setting, leaving its extension to property-guided molecule generation unaddressed.
Conditional molecule generation involves designing new molecules that satisfy specific constraints or target properties. This is typically implemented by concatenating property values with node features to embed property information into molecular graphs, a technique that has been primarily demonstrated on the QM9 dataset [7, 8, 27, 25, 28]. While this approach has proved effective, it may oversimplify the interaction between molecular graphs and properties. For example, Gebauer et al. suggest that applying a Gaussian expansion to a scalar property may help extract property information more robustly [29]. Systematic studies of property embedding methods are lacking, and optimal embedding strategies may vary depending on the specific property. Moreover, previous methods use a pretrained classifier as the property predictor. Since this classifier is trained on the QM9 dataset, where all molecules are fully relaxed, it may struggle to make accurate predictions for generated molecules, which are likely not relaxed. Consequently, the reported results can be unreliable due to the limited generalization capacity of the pretrained classifiers/regressors used for evaluation. Last but not least, property-guided generation is typically evaluated in an ‘in-distribution’ setting, where property values sampled from within the QM9 distribution are used to generate molecules expected to exhibit the corresponding property values. In practice, for molecular discovery, we may want to evaluate the model’s ability to push the envelope of the original training distribution and generate molecules with underrepresented property values.
In this work, we introduce PropMolFlow—a novel property-guided molecule generation framework—that integrates various property embedding methods with an SOTA SE(3)-equivariant flow matching process based on geometric vector perceptrons (GVP). We evaluate PropMolFlow on the QM9 dataset, showing that it outperforms previous diffusion models and flow matching methods, setting a new SOTA for property-guided molecule generation in 3D while achieving much faster inference times. We perform a systematic study of various property embedding methods and identify the optimal approach by property type and property evaluation tasks. We achieve SOTA performance of property-guided generation across many properties while maintaining SOTA performance for standard metrics, such as molecule stability and PoseBusters validity. We propose a new task for property-guided generation that evaluates PropMolFlow models’ capability in generating out-of-distribution molecules. In a first-of-its-kind analysis, we perform extensive density functional theory-based (DFT) calculations to validate the accuracy of our property-guided generation. We highlight the importance of DFT calculations for different tasks and properties, and release a curated DFT dataset of 10,773 molecules generated by PropMolFlow for broader community use. We propose “closed-shell ratio” and revised atomic and molecular stability as key metrics based on valence-electron configuration to filter out open-shell molecules and molecules with invalid valency–charge configurations, because the QM9 training data consists exclusively of closed-shell and charge-neutral molecules.
Results
Overview of PropMolFlow
Figure 1: Overview of the PropMolFlow methodology. a, PropMolFlow models are jointly trained on a molecular graph and property embedding. A molecular graph includes node scalar features, node position features and edge bond features. A property embedding comprises an optional Gaussian expansion mapping (G.E.) followed by a multilayer perceptron (MLP) that projects a scalar property to a high-dimensional embedding space. Conditioning on the property is achieved by the interaction between the property embedding and node scalar features in the molecular graph. b, A joint flow matching process is used for molecule generation by learning different molecule modalities together. c, Five possible interaction types between a property embedding and a molecular graph. Node scalar features $[A,C]$ interact with a property embedding $P=\phi_{\mathrm{prop}}(k)$ via various operations. An MLP transformation $\varphi_{\theta}$ is applied to convert the dimension back to that of the original $[A,C]$, where necessary. $\odot$’ represents an element-wise Hadamard product, and $\oplus$’ stands for a ‘Concatenate’ operation. d, Gaussian expansion as augmented property embedding. Curves in the top panel correspond to five Gaussian basis functions that are evenly spaced between the minimum and maximum gap values. Centers of Gaussian basis functions are marked by gray dashed lines, and the red dashed lines represent the three example input values. The bottom panels show the function values for each Gaussian for two example inputs. In a molecule ball-and-stick representation, white, gray, red, and blue indicate H, C, O, and N, respectively.
PropMolFlow builds on top of the state-of-the-art generative process using flow matching. A flow model learns the molecule graph distribution through training a velocity field conditioned on observed molecule samples. As shown in Figure 1a, PropMolFlow uses a fully-connected graph to represent a molecule and an SE(3)-equivariant graph neural network (GNN) to parameterize the velocity field. Property-guided generation is achieved through the interaction of a property embedding with the node scalar features of a sampled molecular graph. The property embedding is constructed by first transforming the raw property value using a fixed (non-trainable) Gaussian basis expansion, followed by a shallow multilayer perceptron (MLP) that produces the final trainable embedding. This Gaussian expansion is optional, and its utility depends on the specific property type and evaluation metric. Once trained, the SE(3)-equivariant graph neural network defines a joint flow matching process that denoises the molecular graph by simultaneously updating all modalities—such as atom types, bond orders, and 3D coordinates—to generate valid molecules (Figure 1b). Details of the joint flow matching are provided in Methods.
We explored five specific operations, including ‘Concatenation’, ‘Sum’, ‘Multiply’, ‘Concatenate + Sum’, and ‘Concatenate + Multiply’, for combining property embeddings with node scalar features (Figure 1c). The dimension of node scalar features is preserved after interacting with the property embedding, providing a seamless integration into the original architecture. The non-trainable Gaussian expansion layer is used to enrich the property representations prior to being mapped to a property embedding via an MLP. Figure 1d illustrates an example of Gaussian expansion for the HOMO-LUMO gap. The top panel shows the five Gaussian basis functions and the bottom panel lists the function values for two example target properties, indicating the activation of each Gaussian basis function. We investigate the effectiveness of different property embedding strategies for six molecular properties in the QM9 dataset. The property guidance design space includes ten possible methods depending on whether the Gaussian expansion is used and which type of property embedding operation is employed. Details of the Gaussian expansion are provided in Methods.
We trained PropMolFlow on the QM9 data with explicit hydrogen atoms [28, 30]. The QM9 data contains 133,885 molecules, and each molecule contains up to 9 heavy atoms and consists of 3–29 atoms (average: 18 atoms), composed of up to 5 elements: C, N, O, F, and H. Molecules are fully optimized and molecular properties are obtained with DFT at the level of B3LYP/6-31G(2df,p). For property-guided generation, we considered six molecule properties: polarizability ($\alpha$), HOMO-LUMO gap, HOMO energy, LUMO energy, dipole moment ($\mu$), and heat capacity ($C_{v}$). To generate molecular graphs with explicit bond and charge information, we employed the QM9 SDF file distributed by DeepChem [30]. Upon inspection, we discovered numerous inconsistencies: invalid bond orders and non-zero net charges in roughly 30,000 molecules, despite QM9’s requirement for charge-neutral, closed-shell valency [28]. We corrected these discrepancies by reassigning bond orders or charges to enforce valency-charge consistency for the vast majority of entries and used the corrected data to train our PropMolFlow models. After the correction, we improved the structural validity of generated molecules significantly. Previous works that used bond orders overlooked this issue [23, 26]; instead, they modified evaluation metrics to tolerate chemically inconsistent bond–charge combinations, hence inflating molecule stability estimates contrary to QM9’s intended constraints. A detailed comparison between models using previous problematic data and the corrected data, together with the comparison between previous stability metrics and the revised ones, can be found in Supplemental Table LABEL:tab:data_and_metric_compare. A detailed description of our procedure and the alignment with the original QM9 XYZ data is provided in Supplemental Section LABEL:sec:qm9_data_fix. The revised QM9 SDF data is provided on ColabFit Exchange [31, 32].
PropMolFlow achieves state-of-the-art property-guided generation
We evaluated PropMolFlow based on its ability to accurately generate molecules with the target property, the validity of the generated molecules (i.e., whether they adhere to chemical rules), and the speed of generation. These assessments were first conducted under the in-distribution (ID) property-guided molecule generation task. The ID task involves sampling a range of property values from the same distribution as the training data. As such, it assesses the model’s capability to generate structures spanning a broad spectrum of target properties and has been widely adopted in prior studies of property-guided generative models [8, 21, 25, 27].
In this task, a conditional generative model was first used to generate molecules conditioned on property values sampled from the same distribution as the training data (e.g., QM9). These sampled property values served as the input conditions. A separate property predictor/regressor was then applied to estimate the properties of the generated molecules. The predicted property values were compared to the original input conditions, and the mean absolute error (MAE) was used as the evaluation metric. A lower MAE indicates more accurate property conditioning and, therefore, better model performance. In PropMolFlow, we used an SE(3) GVP network for molecule generation and, correspondingly, trained a separate GVP regressor for each of the six molecular properties. For evaluation, we generated 10,000 molecules with input properties sampled from the property distribution of QM9 training data. Further details of the GVP regressors are provided in Methods.
We compared PropMolFlow with five SOTA property-guided generative models, all trained on the QM9 dataset. Among these models, EEGSDE (Equivariant Energy Guided Stochastic Differential Equations) employs an energy function to guide the diffusion for property-guided generation [27]; EquiFM is the first equivariant flow matching model for 3D molecule generation [21]; GeoLDM introduces the first latent diffusion model for molecular geometry generation [8]; Geometry-Complete Diffusion Model (GCDM) constructs an SE(3) equivariant geometry-complete diffusion model using geometry-complete perceptrons for molecule generation [25, 12]; and JODO utilizes a diffusion graph transformer for the joint generation of 2D molecular graphs and 3D geometries. Overall, PropMolFlow achieves competitive performance against the SOTA models on property-guided generation (Table 1). It should be noticed that EEGSDE, EquiFM, GeoLDM, and GCDM do not include bond orders in their molecular graphs while JODO incorporates both bond existence and bond orders. All five baseline conditional models omit atomic charges, in contrast to PropMolFlow, which can generate molecules with non-zero atomic charges (see Methods for more details of the baseline models). The “Random (Upper-bound)” intuitive baseline fully shuffles property values, thereby removing any correlation between molecular structures and properties, and serves as an upper bound on error. The “# Atoms” baseline uses only the number of atoms as a predictor for molecular properties, capturing coarse size-dependent trends. The “QM9 (Lower-bound)” baseline corresponds to a pretrained EGNN regressor provided by Hoogeboom et al. [7], trained directly on QM9 and serving as a lower bound on achievable error. Improvement over the “Random” baseline indicates that the conditional generation effectively integrates property information. Surpassing the “# Atoms” baseline suggests that the generative model captures structural features beyond simple atom count when generating new molecules.
Table 1: Performance of PropMolFlow with respect to property alignment. Results for the baseline models use an equivariant graph neural network (EGNN) as the regressor, whereas PropMolFlow results use our pretrained SE(3) equivariant GVP regressor. Best results are bolded, whereas the second-best results are underlined. JODO results were reported on our sampled molecules using publicly available model checkpoints.
Property $\alpha$ $\Delta\epsilon$ $\epsilon_{\rm{HOMO}}$ $\epsilon_{\rm{LUMO}}$ $\mu$ $C_{v}$
Units Bohr3 meV meV meV Debye cal/(mol $\cdot$ K)
QM9 (Lower-bound) 0.10 64 39 36 0.043 0.040
Random (Upper-bound) 9.01 1470 645 1457 1.616 6.857
# Atoms 3.86 866 426 813 1.053 1.971
EEGSDE 2.62 542 302 496 0.858 1.037
EquiFM 2.41 591 337 530 1.106 1.033
GeoLDM 2.37 587 340 522 1.108 1.025
GCDM 1.97 602 344 479 0.844 0.689
JODO 1.44 333 231 260 0.620 0.580
PropMolFlow 1.31 391 254 315 0.620 0.626
Notably, PropMolFlow shows the lowest MAE for two properties—$\alpha$ and $\mu$—and comparable performance for $C_{v}$ and $\epsilon_{\rm{HOMO}}$ versus the SOTA JODO, and the second best performance for the other two properties $\epsilon_{\rm{LUMO}}$ and $\Delta\epsilon$. For instance, PropMolFlow achieves substantially lower MAEs compared to GCDM for $\epsilon_{\rm{LUMO}}$ (315 meV vs. 479 meV), and $\mu$ (0.620 Debye vs. 0.844 Debye), demonstrating a clear advantage in controllable molecular generation over prior diffusion- and flow-based approaches. The superior performance of PropMolFlow can be ascribed to the inclusion of bond order information and discrete treatment for atom types, charges and bond orders, which encodes structural information more accurately and hence offers a more robust structure–property mapping for generated molecules. It also benefits from the thorough exploration of property-dependent optimal embedding methods (Supplementary Table LABEL:tab:mae_range and Supplemental Section LABEL:sec:model_stats). For instance, the MAE variation for $\Delta\epsilon$ can vary by more than 40%, from 391 meV using a ‘Concatenate_Sum’ embedding to 551 meV using a ‘Concatenate’ embedding, both without Gaussian expansions.
Rapid molecule generation with high structural fidelity
Figure 2: Chemical validity and sampling efficiency of PropMolFlow against five baseline models. a, Molecule stability. b, RDKit validity. c, Uniqueness. d, PoseBusters validity. e, Closed-shell ratio. f, Sampling time for 10k molecules. Mean $\pm$ standard deviations are reported across six molecule properties. Structural validity metrics are shown in the value range between 0.4 and 1, while sampling times are in the unit of minutes.
Besides the ability to generate molecules with target properties, the structural validity and speed of generation are also critical in ensuring the success and robustness of chemical discovery with generative models. We further evaluated the performance of PropMolFlow by assessing the structural validity of the generated molecules and the sampling efficiency. Structural validity refers to whether the generated molecules conform to basic chemical rules. This was measured using several metrics, including atomic stability, molecule stability, RDKit validity, ‘uniqueness and validity’, PoseBusters validity, and closed-shell ratio. Atom stability is given by the proportion of atoms with correct valency. To account for atomic charges, an atom is considered to have the correct valency if its formal charge balances its explicit valency, as defined in RDKit [33]. For instance, a nitrogen atom should carry a +1 formal charge if it has a valency of 4. Molecule stability defines the proportion of molecules in which all atoms are stable and the molecule is charge neutral (zero net formal charges). RDKit validity refers to the ratios of molecules that pass RDKit’s sanitization check. ‘Uniqueness and validity’ is defined as the proportion of molecules that are RDKit valid and unique in their SMILES representation. PoseBusters validity corresponds to the fraction of molecules that pass many de novo chemical and structural validation tests, such as all atom connectivity, valid bond lengths and angles, and absence of internal clashes [34]. A molecule is closed-shell valid if it has an even number of valence electrons. Since all molecules in QM9 are closed-shell, any generated open-shell species are considered invalid. Sampling efficiency is given by the wall-clock time required to generate 10,000 molecules. Computational settings for sampling are provided in Methods.
Across nearly all stability and validity benchmarks, PropMolFlow consistently outperforms baseline modes (Figure 2a–e). The only exception is the closed-shell ratio, where it slightly underperforms JODO—a diffusion model that omits atomic formal charges and adds additional bond features such as aromaticity existence. Moreover, PropMolFlow achieves high chemical validity with much faster sampling efficiency. Flow matching is normally much faster than diffusion models due to its shorter, deterministic probability path (Eq. (6) and (8)) and optimal transport (Supplementary Section LABEL:sec:ot[35, 21]. PropMolFlow requires only 100 time steps compared to the 1000 diffusion time steps, and achieves a speedup for sampling of at least 8$\times$ over diffusion-based models and it is nearly 2$\times$ faster than EquiFM (Figure 2f).
When evaluating metrics on a per-property basis, PropMolFlow consistently achieves the highest molecular stability and validity across all properties (Supplementary Tables LABEL:tab:propmolflow_metric_by_property-LABEL:tab:jodo_metric_by_property). Across properties, PropMolFlow’s metrics follow the order: $\mu\approx\Delta\epsilon>\epsilon_{\mathrm{HOMO}}>\epsilon_{\mathrm{LUMO}}>\alpha>C_{v}$. This trend generally holds for the baseline models as well, with the exception of EEGSDE, suggesting a more complex structure–property relationship for $\alpha$ and $C_{v}$. This complexity may be due to their strong dependence on molecular geometry and the limitation of the training data, which includes only fully relaxed molecules. Across metrics, molecular stability and “Validity & Uniqueness” exhibit over 2% variations across properties, whereas molecular validity and PoseBusters validity vary by less than 1%. Besides, all methods in this study generate a non-negligible fraction of open-shell molecules: on average 1.6% for JODO and at least 4% for the other methods. This highlights a previously overlooked issue in generative molecular modeling: the inability to consistently respect the valence electron constraints inherent to the training data. Addressing this limitation is essential for improving the chemical validity of generated molecules especially when atomic formal charges are considered in molecular graphs, and future work should improve upon this direction—perhaps by explicitly including closed-shell constraints, as we do here by adding bond orders into molecular graphs—to ensure better alignment with the reference distribution.
Property predictor as DFT proxy reveals systematic inductive bias
Figure 3: Performance of GVP property predictors for PropMolFlow-generated molecules without and with relaxation. a, Comparison between Target, DFT and GVP shows the reliability of GVP to evaluate the MAE metrics commonly used for property-guided generation. Comparison between GVP and GVP-R, and between DFT and DFT-R shows the structural dependence of GVP-predicted and DFT-predicted properties, respectively. Comparison between GVP and DFT with and without relaxation shows the reliability of GVP in capturing the ground-truth DFT values over raw and relaxed structures. Target is the input target property values used by PropMolFlow models to generate models. ‘DFT’ and ‘GVP’ are respective DFT-computed and GVP-evaluated property values on directly generated raw molecules. ‘DFT-R’ and ‘GVP-R’ refer to values evaluated on DFT-relaxed molecules. b, Pairwise comparison between Target, DFT and GVP on raw molecules. c, GVP versus DFT for both raw and DFT-relaxed molecules. d, Structure variation and property dependence on the DFT relaxation. Molecules with the highest RMSD for each property and their DFT and GVP values are listed as examples. In molecule representation, gray, red, blue, and white colors indicate C, O, N, and H, respectively. Property values for $\alpha$, $\Delta\epsilon$, and $C_{v}$ are in units of Bohr3, eV, and cal/(mol $\cdot$ K), respectively.
A separate surrogate model is needed to evaluate the properties of generated molecules. Typically, this surrogate model shares the same architecture as the generative neural network model [7]. However, using models from the same family may introduce inductive biases, making it necessary to validate results using external property predictors, such as DFT. To this end, we performed DFT calculations on 1000 molecules selected from the 10,000 structures generated for each property in the ID task. For directly generated molecules, multi-step filtering is essential to ensure that the structures are chemically valid and suitable for downstream applications. We filtered the 1000 molecules by five criteria: Molecule stability, RDKit validity, PoseBusters validity, closed-shell validity, and multi-fragment check. The multi-fragment check filters out molecules that with more than one disconnected fragment. For molecules passing all filters, property values are evaluated using both a GVP property predictor and DFT calculations. We also filtered out molecules whose DFT calculations did not converge, resulting in 90% of the molecules remaining for the subsequent analyses (Table LABEL:tab:no_sp_dft). To evaluate the reliability of GVP property predictor and the impact of structural relaxation, we conducted a comprehensive comparison among target property values (Target), DFT-calculated properties (DFT), and GVP predictions (GVP)—using both directly generated and DFT-relaxed molecules (Figure 3a). We chose three example properties, $\alpha$, $\Delta\epsilon$ and $C_{v}$, for the following analyses because the results for the other three properties are similar to those of $\Delta\epsilon$ (Supplemental Figure LABEL:fig:si_target_dft_gvp). Example orders of pairwise MAE distances for $\alpha$ are also provided in Figure 3a.
We first assess the reliability of GVP on the initially generated molecules (Figure 3b). The MAEs between the target values and DFT (Target vs. DFT) are comparable to those betwen the target and GVP predictions (Target vs. GVP) for $\alpha$ and $\Delta\epsilon$, suggesting that GVP property predictors are statistically reliable for estimating properties of directly generated molecules. However, GVP consistently underestimates the MAEs, revealing an inductive bias in its predictions. Notably, although the MAEs of GVP and DFT relative to the target are similar, the MAEs between GVP and DFT themselves are significantly larger—for example, 0.91 vs. 0.1 Bohr3 for $\alpha$, and 204 vs. 43 meV for $\Delta\epsilon$. For $C_{v}$, the MAE between the target values and DFT is much larger than that between the target values and GVP. However, this discrepancy is not due to poor reliability of the GVP property predictors, but rather stems from an issue with DFT-calculated heat capacities ($C_{v}$) on unrelaxed structures. DFT-computed $C_{v}$ values depend on vibrational frequencies, which are highly sensitive to molecular geometry. To validate this analysis, we relaxed the structures, resulting in the DFT vs. Target MAE dropping from 1.68 to 0.68 cal/(mol $\cdot$ K), closely aligning with the MAE of 0.61 cal/(mol $\cdot$ K), predicted by GVP.
Evaluation on directly generated molecules without geometry optimization reflects the statistical quality of the learned structure–property relationships captured by both PropMolFlow models and the GVP predictors. However, from a chemical discovery perspective, it is essential to relax the structures and evaluate properties based on the relaxed chemical stable conformations that are more likely to occur in real-world experiments. To assess the reliability of GVP on relaxed structures, we compare its predictions with DFT-computed property values for both unrelaxed and relaxed molecules (Figure 3c). For $C_{v}$, a much closer agreement between DFT and GVP was observed on relaxed structures. In contrast, for $\alpha$, a notable offset between GVP and DFT appears when using relaxed structures, indicating that GVP is more reliable on unrelaxed structures. This can be attributed to the inductive bias of GVP property predictors—specifically, GVP tends to perform better on directly generated structures by PropMolFlow, as both models share the same architecture. This bias is further supported by the increased MAEs between GVP and DFT after relaxation for all properties except $C_{v}$ (Supplemental Figure LABEL:fig:id_mae_heatmap). For $\Delta\epsilon$, despite a slight broadening of the peak at lower $\Delta\epsilon$ values is observed, both DFT and GVP predictions remain largely unchanged, suggesting that this property has weak dependence on structural relaxation.
To further quantify the effect of relaxation, we calculated the symmetry-aware root mean square distances (RMSDs) between atomic positions of raw structures and DFT-relaxed structures. Given the same chemical compositions and element-type ordering, the RMSD between any two molecules (M0 and M1) is the minimum distance between these two structures considering the translational and rotational symmetry operations, which can be expressed as:
$\mathrm{RMSD}(\rm{M}_{0},\rm{M}_{1})=\min_{P\in SE(3)}\{d\left(M_{0},P(M_{1})\right)\}$ (1)
Where $SE(3)$ represents the group of operations that respect the translational and rotational symmetry. The RMSD distributions for $\alpha$, $\Delta\epsilon$ and $C_{v}$ are showed in Figure 3d. All RMSD distributions are highly left-skewed and concentrated around 0 Å, meaning the majority of PropMolFlow-generated structures are close to their relaxed, stable counterparts. To estimate the property sensitivity to structural relaxation, we selected the structures whose RMSDs are larger than 0.03 Å—RMSDs lower than this threshold are considered very close to the relaxed structures, and in Eq. (2), we define the normalized structure sensitivity of DFT-calculated properties to allow comparisons between different properties.
$\chi_{\rm{DFT}}=\frac{|{\delta q_{\rm{DFT}}|}}{(q_{0.99}-q_{0.01})\cdot\rm{RMSD}}$ (2)
where $\delta q_{\rm{DFT}}$ is the change of DFT-calculated property upon relaxation, and the normalizer ($q_{0.99}-q_{0.01}$) is the difference between the 99th quantile and the 1th quantile of the property distribution of QM9 training data. This range is chosen over the min-max range because it is robust to outliers. Despite similar RMSD distributions across the three properties, the differences are more pronounced in the structural dependence of DFT-calculated properties, following the order of $C_{v}>\alpha>\Delta\epsilon$. On the right panel of Figure 3d, molecule configurations with the highest RMSDs are shown, and their RMSD, taget input, DFT-calculated, GVP-predicted property values are listed to rationalize the relationship between GVP, DFT and Target values with and without relaxation. DFT property values for relaxed structures can be either closer or farther to the target input values. For these molecules with large structural changes, the GVP predictions are nearly unaltered, showing a much weaker structural dependence compared to that of DFT. Although DFT property changes are more significant than GVP upon relaxation for the three extreme cases, GVP predictions are close to DFT-calculated values for both raw and relaxed molecules, offering reliable evaluation of PropMolFlow models and robust guidance towards chemical discovery.
Interpolated structures follow chemically intuitive trends
To assess whether PropMolFlow models have learned a continuous and smooth structure–property relationship, we performed interpolation by varying the target property values while fixing the number of atoms in the generated molecules. For each property value, ten molecules were sampled, and the one that was chemically valid and whose DFT-calculated property most closely matched the target property value was selected. For each property, we show the minimum and maximum target values along with their DFT-calculated values below the molecular configurations. The target property value, DFT-calculated value, and GVP predictions for each configuration are provided in Supplemental Table LABEL:tab:interpolation.
Figure 4: PropMolFlow generated molecules when interpolating among different property values while fixing the number of atoms at 19 for sampling. The minimum and maximum target properties in red and corresponding DFT-calculated properties in blue are provided below the configurations. All molecules chosen here pass the five filtering criteria and have DFT values closest to the target properties. C, H, O, and N are in gray, white, red, and blue colors, respectively. Property units are provided in the square brackets under each property symbol.
[…]
The interpolated structures generated by PropMolFlow exhibit smooth and chemically intuitive transitions as property values vary, demonstrating that the model has learned meaningful structure—property relationships. These results highlight the model’s ability to generate property-guided molecules that align with established chemical principles.
Figure 5: Performance of PropMolFlow in OOD generation. a, Distribution of DFT calculated and GVP predicted values for PropMolFlow generated molecules, and the property distribution of QM9 is also included. The vertical black dashed line in histograms represents the target property value $q_{0.99}$. Curves on top of histograms are fitted with a kernel density estimation. b, Three example molecules that do not exist in QM9 but are found in a larger PubChem dataset are included in the left panel. Numbers below the configurations are DFT calculated property values on raw molecules generated by PropMolFlow models. C, H, O, N, and F are in gray, white, red, blue, and yellow colors, respectively. Property values for $\alpha$, $\Delta\epsilon$, and $C_{v}$ are in units of Bohr3, eV, and cal/(mol $\cdot$ K), respectively.
PropMolFlow pushes molecule generation toward out-of-distribution property space
Although ID task is typically used to evaluate property-guided generative models, it is also of practical interest to test molecular generation for properties that are underrepresented in the training dataset. To this end, we propose an ‘out-of-distribution’ (OOD) task designed to assess the model’s ability to generate molecules conditioned on property values that are rare or sparsely represented in the original training distribution. In this task, we chose the 99th quantile of the property distribution of QM9 training data as the target property value. The target value for each property is provided in Supplemental Table LABEL:tab:prop_choice. We sampled 1,000 structures conditioned on the chosen property value using the models with top-performing embedding methods (Supplemental Table LABEL:tab:best_model). These generated molecules then pass through the same multi-step filtering procedure used in the ID task, and their properties are evaluated by both GVP and DFT.
Here we compare the property distribution of generated molecules against that of the QM9 training data for $\alpha$, $\Delta\epsilon$ and $C_{v}$ (Figure 5a). Results for other properties are provided in Supplemental Figure LABEL:fig:si_ood. The DFT results demonstrate that the OOD generation successfully shifts the property distribution of generated molecules towards the specified target values (indicated by the dashed line), and the DFT-predicted distribution is nearly perfectly centered on the target property value for the polarizability ($\alpha$). However, for $\Delta\epsilon$ and $C_{v}$, the DFT-predicted distribution is off the target by a small margin. For example, both the GVP and DFT distributions for the gap are centered around 8.95 eV, which is 0.41 eV away from the target value of 9.36 eV. This deviation is likely due to the scarcity of training data in the corresponding region of property space. One should note the DFT issue for $C_{v}$ on unrelaxed structures which can be largely addressed by performing relaxation. While property guidance effectively steers the generative model toward the desired values, the model’s capacity to extrapolate is ultimately constrained by the coverage of the training distribution. Other guidance approaches, such as classifier-free guidance [36], can adjust the weights for the property conditioning, potentially achieving better alignment with target property values, as demonstrated in the conditional generation of inorganic crystal materials [37]. Moreover, the GVP distributions are almost overlapped with the DFT ones, showing desired extrapolation capability of the GVP property predictors. To further evaluate the novelty of generated molecules, Figure 5b presents three example molecules that are not part of the QM9 dataset but are found in the larger PubChem database [38]. These molecules exhibit the DFT properties close to their target input values. This demonstrates PropMolFlow’s ability to generate chemically valid molecules with desired target properties that lie outside the original training domain.
Conclusion
In this work, we introduce PropMolFlow, a property-guided molecule generation framework based on the SOTA SE(3) flow matching. Our approach outperforms existing diffusion and flow matching models in the conditional generation setting, offering improved property alignment, structural validity and sampling efficiency across many molecular properties. We conduct a systematic evaluation of property embedding strategies and identify optimal configurations that depend on both the target property and the task regime—whether ID or OOD. To enable a more comprehensive evaluation, we propose two new metrics and one new task: The closed-shell ratio and revised atomic stability and molecular stability to account for charged species, which quantify how well the generated molecules align with the training data in terms of valency–charge configurations, and an OOD task, designed to evaluate controllability in sparsely represented regions of property space.
To validate the fidelity of generated molecules and reliability of property predictors, we perform extensive DFT calculations. PropMolFlow generates molecules that are close to the stable, relaxed counterparts. Generally, GVP property predictors provide reliable approximations of the true DFT MAEs against target values despite the relative large difference between GVP and DFT. We demonstrated the inductive bias of property predictors, which lead to a systematic underestimation of true MAEs and its partial failure when predicting on unseen relaxed structures. Additionally, we curated a DFT dataset of generated molecules, which is made available through the ColabFit Exchange [31, 39]. While this work focuses on small molecules from the QM9 dataset, PropMolFlow is readily extendable to larger molecular datasets with more atoms, as well as to materials datasets, as long as property labels are available.
Methods
Joint Flow Matching for Molecule Generation
Ideally, new molecules would be generated by sampling directly from the probability distribution over valid molecular graphs $q(x)$. However, this target distribution lacks a tractable analytical form and can only be approximated through a generative process. To this end, we adopt flow matching to learn an approximation of $q(x)$. This flow matching is parameterized by an SE(3) equivariant graph neural network built on top of FlowMol architecture [26], details of which are provided in Supplemental Section LABEL:sec:appendix_model. In flow matching, the objective is to learn a time-dependent velocity field $u_{t}^{\theta}$ parameterized by a neural network $\theta$, which defines a probability path $p_{t}(x):[0,1]\times x_{0}\rightarrow x_{1}$ from a base distribution, $p_{0}(x)$ to a target distribution, $p_{1}(x)=q(x)$.iiiWe use subscripts to denote time dependence, e.g., $p_{t}(x)$ The flow induced by the velocity field is given by the ODE $\dot{x}=u_{t}\left(x\right)$, which describes the evolution of samples along the probability paths. Designing a flow matching process thus involves two key steps: 1) specifying a probability path that satisfies the boundary conditions $x_{0}\sim p_{0}$ and $x_{1}\sim p_{1}$, and 2) training a neural network $\theta$ to approximate the velocity field that transports samples along this path.
In practice, while accurately modeling the exact velocity field is challenging, conditioning it on a random variable $z$ leads to a more tractable formulation. Importantly, optimizing the conditional velocity field is equivalent to optimizing the marginal velocity field [40]. The endpoint reformulation of [24] reparametrizes the flow matching problem by focusing on learning a denoising function rather than the full time-dependent velocity field $u_{t}^{\theta}$. Specifically, the method learns a denoising function $x^{\theta}_{1|t}$ that predicts the final state $x_{1}$ from the intermediate state $x_{t}$. The training objective then minimizes the discrepancy between the predicted final sample and the true final sample:
$\mathcal{L}=\mathbb{E}_{t,p_{t|1}(x_{t}|z),p_{z}}\left[\lVert x^{\theta}_{1|t}-x_{1}\rVert\right],$ (3)
where $z$ is a conditioning variable that can be either the final graph $x_{1}$ or a pair $z=(x_{0},x_{1})$ representing both the initial and final states.
Discrete graph features are handled via CTMC as developed by Campell et al. [20] and Gat et al. [41]. In this setting, Eq. (3) is reformulated to minimize a cross-entropy loss over discrete variables:
$\mathcal{L}_{\rm{CE}}=\mathbb{E}_{t,p_{t|1}(x_{t}|z),p_{z}}\left[-\log p^{\theta}_{1|t}(x_{1}^{i}|x_{t})\right]$ (4)
We factorize the conditional probability path of the molecular graph into the individual paths for each modality. A joint flow matching process is then defined by learning a velocity field (or, equivalently, a denoiser)—parameterized by a GNN—that minimizes the total loss expressed as a weighted sum of the losses for each modality:
$\mathcal{L}=\eta_{X}\mathcal{L}_{X}+\eta_{A}\mathcal{L}_{A}+\eta_{C}\mathcal{L}_{C}+\eta_{E}\mathcal{L}_{E}$ (5)
Previous studies suggest that atomic positions should have a higher weight, followed by bond orders, atomic charges, and atom types. The loss weights are chosen to be $(\eta_{X},\eta_{A},\eta_{C},\eta_{E})=(3.0,0.4,1.0,2.0)$. The factorization of the flow into its components allows for the flexible design of individual interpolants for each modality, as detailed below.
Interpolant for atomic positions.
The atomic position flow is conditioned on the pair of initial and final states $z_{X}=(X_{0},X_{1})$, and is given by the linear interpolant
$X_{t}=\alpha_{t}X_{0}+\beta_{t}X_{1}$ (6)
where $\alpha_{0}=\beta_{1}=1$ and $\alpha_{1}=\beta_{0}=0$. The functions $\alpha_{t}$ and $\beta_{t}$, known as the interpolant schedule, control the rate of mixing between the base and target distributions. A typical and effective choice of interpolant schedule for both molecule and materials generation is a linear schedule: $\alpha_{t}=(1-t)$ and $\beta_{t}=t$ [26, 18]. Other variants of interpolants are discussed in detail in [17], and they have been validated in the context of inorganic crystals generation in our OMatG framework [18]. The base distribution for the atomic position flow is a standard Gaussian distribution $p_{0}(X)=\prod_{i=1}^{N}\mathcal{N}(X_{0}^{i}|\mathbf{0},\mathbb{I}_{3})$. A denoising network is then trained to produce the final atomic positions $X^{\theta}_{1|t}$, minimizing the following loss:
$\mathcal{L}_{X}=\mathbb{E}_{t,p_{t}(X_{t}|X_{0},X_{1}),\pi(X_{0},X_{1})}\left[\lVert X^{\theta}_{1|t}-X_{1}\rVert\right]$ (7)
Where the joint distribution $\pi(X_{0},X_{1})$ defines the optimal transport coupling between $(X_{0},X_{1})$. Details of the optimal transport formulation are provided in Supplementary Section LABEL:sec:ot.
Interpolant for atom type, charge, and bond order.
Atom types, charge and bond orders are modeled using CTMC-based flows [20]. For each categorical variable, a ‘mask’ token is added as an additional discrete state. For example, consider atomic identities $A=\{A^{1},A^{2},\ldots,A^{N}\}$ where each $A^{i}$ takes values in the set $\{1,2,\ldots,n_{A},M\}$. Here, $n_{A}$ is the number of predefined atom types and $M$ is the mask token. The joint probability path over all atomic identities is factorized into independent contributions from each atom. The conditional probability path for atom $i$ is defined as:
$p_{t}(A^{i}_{t}|A_{0},A_{1})=\alpha_{t}\delta(A^{i}_{t},A_{1}^{i})+\beta_{t}\delta(A^{i}_{t},M)$ (8)
where $\alpha_{t}$ and $\beta_{t}$ are the same interpolant schedules defined in Eq. (6). A linear interpolant is effective for categorical variables as well [26, 18]. Here, $\delta(i,j)$ denotes the Kronecker delta, which is 1 if $i=j$ and 0 otherwise. The base distribution corresponds to all atoms being initialized in the masked state. This formulation implies that at time $t$, the atom identity $A^{i}_{t}$ has a probability $\alpha_{t}$ of being in its final state $A_{1}^{i}$ and a probability $\beta_{t}$ of being in the masked state $M$.
In contrast to continuous variables, discrete flows are propagated via CTMC, where state transitions are stochastic rather than deterministic. The transition probability for atom $i$ over a short time interval $\Delta t$ can be expressed in terms of individual atomic contributions:
$p^{i}_{t+\Delta t}(j|A^{i}_{t})$ =$=\left\{\begin{array}[]{lcl}R^{i}(A^{i}_{t},j)&\mbox{for}&j\neq A_{t}^{i}\\ 1+R^{i}(A^{i}_{t},A_{t}^{i})&\mbox{for}&j=A_{t}^{i}\end{array}\right.$ (11)
$=\delta(j,A_{t}^{i})+R^{i}(A^{i}_{t},j)\Delta t$ (12)
where $R(A^{i}_{t},j)$ is the rate matrix specifying the transition probability from atom type $A^{i}_{t}$ to $j$ [20]. To ensure proper normalization of the transition probabilities at time $t+\Delta t$, the rate matrix must satisfy $R^{i}(A^{i}_{t},A^{i}_{t})=-\sum_{j\neq A^{i}_{t}}R(A^{i}_{t},j)$ so that the total probability mass remains normalized.
An additional stochastic term $\eta$ can be introduced into the rate matrix to form the modified rate matrix $R_{t}^{\eta}:=R_{t}^{*}+\eta R_{t}^{\mathrm{DB}}$. This new rate matrix generates the same marginal probability path if $R_{t}^{\mathrm{DB}}$ satisfies detailed balance over the conditional probability path $p_{t|1}$. By including the stochasticity term and choosing a linear interpolant ($\alpha_{t}=1-t$ and $\beta_{t}=t$), the rate matrix for atom $i$ becomes:
$R^{i}(A_{t}^{i},j)=\frac{1+\eta t}{1-t}p^{\theta}_{1|t}(j|A_{t}^{i})\delta(A_{t}^{i},M)+\eta\left(1-\delta(A_{t}^{i},M)\right)\delta(j,M)$ (13)
This formulation implies that if atom $i$ is in the masked state at time $t$, the probability of transitioning to atom type $j$ at time step $t+\Delta t$ is $\Delta t\frac{1+\eta t}{1-t}$. Conversely, if atom $i$ is in an unmasked state, the probability of transitioning back to a masked state is $\eta\Delta t$. In Eq. (13), the rate matrix can diverge as $t\to 1$. To address this issue, we employ a time-dependent loss function that avoids division-by-zero [42].
Additionally, low-temperature sampling—which rescales the prediction logits to sharpen the output distribution—is found to be critical for improving model performance. Specifically, the denoiser network’s output $p^{\theta}_{1|t}(A_{1}^{i}|A_{t})$ is transformed as follows:
$\widehat{p}^{\theta}_{1|t}(A_{1}^{i}|A_{t})=\mathrm{softmax}\left(\nu^{-1}\log p^{\theta}_{1|t}(A_{1}^{i}|A_{t})\right)$ (14)
where $\nu$ is a hyperparameter. The modified probabilities $\widehat{p}^{\theta}_{1|t}(A_{1}^{i}|A_{t})$ are then employed in the discrete loss function defined in Eq. (4). When training PropMolFlow, we set $\eta=10$ and $\nu=0.05$. During inference, we observe that turning off the stochastic term leads to better performance. To ensure that the generated molecules satisfy the required symmetry constraints, the joint flow matching process is constructed with an invariant base distribution and an equivariant transition probability path. A formal derivation of the equivariance properties of the flow matching process is provided in Supplementary Section LABEL:sec:equivariantfm.
Property Embedding Methods
Conditional generation can be framed as sampling a property-conditional target distribution $p_{1}(x|k)$ over molecular graphs, where $k$ is a property. Previous approaches to conditional generation concatenate a property value to the node features of a molecular graph, training the model jointly on the property and the graph [7, 8, 21]. As a result, the learned denoising process is conditioned on the property. This approach requires modifying the input dimension of the denoiser to include the additional property value, which is subsequently discarded in the final output. However, directly appending the property to the node features may oversimplify the interaction between property values and molecular graphs. As shown in Figure 1, our implementation maps a scalar property, $k$, to a high-dimensional property embedding space $P=\phi_{\mathrm{prop}}(k)$ using a MLP, $\phi_{\mathrm{prop}}$. The dimensionality of the property embedding matches that of the node scalar features, allowing seamless integration without requiring additional dimensional matching. To encode interactions between the property embedding and node features, we propose five distinct operations. Depending on the operation, if the combined feature vector $[A,C]$ maintains its original dimensionality after interacting with the property embedding, no further processing is applied; otherwise, an MLP is used to project it back to the original dimensionality. Since the property embedding interacts exclusively with node scalar features, leaving position vectors unchanged, the resulting property-conditioned graph retains SE(3) equivariance. Moreover, the loss function preserves its original form, and the trained SE(3) GNN defines a valid conditional generative process.
In the ‘Sum’ operation, the property embedding is simply added to the node scalar features. The ‘Multiply’ operation performs an element-wise Hadamard product $\odot$’. To ensure the multiplicative factors remain within a bounded range, the property embedding is passed through a sigmoid function and then shifted by 0.5, resulting in multipliers in the interval $[0.5,1.5]$. Both ‘Sum and Multiply’ preserve the original feature dimensionality. In contrast, operations involving ‘Concatenate’ increase the dimensionality of the node scalar features, and an MLP transformation $\varphi_{\theta}$ is required to project the combined features back to the original dimension. For the composite operations, ‘Concatenate + Sum’ and ‘Concatenate + Multiply’, the property embedding is first concatenated to the node features, followed by the respective arithmetic operation. In both cases, the final output has the desired dimensionality, either by design or via the subsequent MLP transformation.
Gaussian Expansion for Augmented Property Embedding
We incorporate a non-trainable Gaussian expansion layer to enrich the property representations prior to being mapped to a property embedding via an MLP $\phi_{\mathrm{GE}}$ [29]. The Gaussian expansion maps a scalar property value into a fixed-length vector using a set of Gaussian basis functions. Given a property $k$ with value $\tau_{k}$, the expanded representation takes the form:
$f_{\mathrm{k}}=\phi_{\mathrm{GE}}\left(\left[\exp\left(-\frac{(\tau_{\mathrm{k}}-(\tau_{\mathrm{min}}+n_{g}d))^{2}}{2d^{2}}\right)\right]_{0\leq n_{g}\leq\frac{\tau_{\mathrm{max}}-\tau_{\mathrm{min}}}{d}}\right)$ (15)
where $\tau_{\mathrm{min}}$ and $\tau_{\mathrm{max}}$ are the minimum and maximum property values, $d$ is the uniform spacing between the Gaussian centers, and $n_{g}$ indexes the basis functions. The output vector encodes the proximity of the input $\tau_{k}$ to each Gaussian center, producing a smooth, localized representation that is refined by the MLP $\phi_{\mathrm{GE}}$.
Figure 1d shows an example of a Gaussian expansion for the HOMO-LUMO gap. In this example, five Gaussian basis functions ($n_{g}=5$) are used, with centers evenly spaced between the minimum and maximum gap values in the QM9 dataset: $\tau_{\mathrm{min}}=0.0246$ and $\tau_{\mathrm{max}}=0.6221$ Hartree. The width of each Gaussian is set by the spacing parameter $d=(\tau_{\mathrm{max}}-\tau_{\mathrm{min}})/(n_{g}-1)$. The top panel of Figure 1d shows the five Gaussian basis functions, with their centers marked by vertical dashed gray lines. In the bottom panel, we pick three example property values, namely 0.1, 0.3, and 0.5 Hartree, and show the expanded feature vectors as bar plots indicating the activation of each Gaussian basis function. For instance, if the gap is 0.1 Hartree, the first two Gaussian functions (centered closest to 0.1) have higher responses than the remaining functions, whose centers are farther away. Similar activation patterns are observed for the other two property values.
Datasets and Baselines
Datasets.
We used RDKit’s sanitization function [33] to filter chemically valid molecules on the QM9 SDF file with bond and charge fix, resulting in a curated set of 133k molecules. After random shuffling, we split the dataset into 100k training, 20k validation, and 13k test samples. For property-guided generation, we considered six molecule properties: polarizability ($\alpha$), HOMO-LUMO gap, HOMO energy, LUMO energy, dipole moment ($\mu$), and heat capacity ($C_{v}$). Precise definitions of each property can be found in Supplementary Section LABEL:sec:more_details. We did not evaluate PropMolFlow on the GEOM-Drugs dataset, which is commonly used for unconditional generation and includes larger molecules, because it only provides quantum-mechanical energy values and lacks other relevant properties [43]. PropMolFlow models and GVP property predictors are trained on disjoint 50k and 50k datasets from the 100k training data.
Baselines.
We compare PropMolFlow against several recent property-guided molecular generation methods based on diffusion models and flow matching, including EEGSDE, GCDM, GeoLDM, JODO, and EquiFM [27, 25, 8, 44, 21]. In this work, we focus exclusively on conditional generation results; unconditional generation performance is reported in earlier studies. Language-model-based approaches are excluded from this comparison, as state-of-the-art models in that category rely on fundamentally different molecule representations and training data regimes. Our comparison includes both structural-validity and property-specific evaluation metrics. For EEGSDE, we benchmark PropMolFlow against the canonical conditional model with a scaling factor of 1 ($s=1$[27]. Although EEGSDE provides DFT-based validation for five out of six properties, it does so on only 100 generated molecules, which limits the statistical significance of those results.
EEGSDE reported the atomic and molecular stability results for its conditional generation, JODO reported molecular stability, and GCDM reported PoseBusters validity for the conditional generation of GCDM and GeoLDM models—no other baseline conditional models provided these metrics. Because EEGSDE and GCDM released their model checkpoints, we sampled from those and computed all structural validity metrics on the resulting structures; for GeoLDM, we used the checkpoints bundled with GCDM. EquiFM checkpoints were unavailable, so we trained its conditional models ourselves following information in [21] (training details in Supplementary Section LABEL:sec:more_details).
Except for the closed-shell ratio, all structural validity metrics depend on bond-order information, which baseline models (other than JODO) do not output. We therefore assigned bond orders using the distance-based cutoffs introduced by Hoogeboom [7]. Although recent work [23] suggests that the bond orders can probably be optimized through the OpenBabel program [45]. We did not apply such refinements to the baseline samples.
For property metrics, PropMolFlow is directly compared to prior methods on the ID tasks. OOD results are not directly comparable due to a lack of baseline generative models and property prediction models testing this regime, but we validate our results with DFT calculations to assess physical fidelity.
Computational Settings
Hyperparameters, Training and Inference Details
PropMolFlow uses the model architecture detailed in Supplemental Section LABEL:sec:appendix_model. PropMolFlow models on the QM9 dataset were trained with 8 molecule update blocks. Atoms contain 256 hidden scalar features and 16 hidden vector features. Edges contain 128 hidden features. QM9 PropMolFlow models were trained with 2000 epochs. Training and inference for PropMolFlow models and all baseline models used a single NVIDIA A100-SXM4 graphic card with 80GB of memory, with a batch size of 128. PropMolFlow models can be trained in about 4–5 days, and EquiFM models were trained with 2500 epochs, which took around 3.5 days.
During inference, PropMolFlow models use Euler’s method with 100 evenly spaced time steps to integrate the learned velocity field. The number of atoms for each generated molecule is first sampled from the QM9 atom count distribution, and the property values are then sampled according to the number of atoms to ensure consistency with the dataset’s statistics, as used in previous property-conditioning works [7, 8, 25].
GVP regressor details.
To be self-consistent, we trained property regressors using Graph Neural Networks based on GVPs. GVP property predictors are trained on a disjoint dataset versus that used for training PropMolFlow models, hence avoiding data leakage between the generative models and property prediction models. We appended an MLP layer that takes the final node scalar features as input to predict the target property. The parameters of GVP regressors are optimized by minimizing a mean squared error loss function. Separate models were trained for each of the six tested properties.
Acknowledgment
The authors acknowledge funding from NSF Grant OAC-2311632 and from the AI and Complex Computational Research Award of the University of Florida. S.M also acknowledges support from the Simons Center for Computational Physical Chemistry (Simons Foundation grant 839534). The authors also gracefully acknowledge UFIT Research Computing for computational resources and consultation, as well as the NVIDIA AI Technology Center at UF. We would like to thank Alex Morehead (University of Missouri) for providing the checkpoint models for the conditional generation of GeoLDM.
Data and code availability
Code and data of this work can be found on github: https://github.com/Liu-Group-UF/PropMolFlow.

### v4

@@
 However, diffusion models still lead in property-guided generation.
-In this work, we introduce PropMolFlow, a novel approach for property-guided molecule generation based on geometry-complete SE(3)-equivariant flow matching.
-Integrating five different property embedding methods with a Gaussian expansion of scalar properties, PropMolFlow achieves competitive performance against previous SOTA diffusion models in conditional molecule generation across various properties while preserving the stability and validity of the generated molecules, consistent with its unconditional counterpart.
-Additionally, it enables faster sampling speed with significantly fewer time steps compared to baseline models.
-We highlight the importance of validating the properties of generated molecules through DFT calculations performed at the same level of theory as the training data.
-Specifically, our analysis identifies properties that require DFT validation and others where a pretrained SE(3) geometric vector perceptron regressors provide sufficiently accurate predictions on generated molecules.
-Furthermore, we introduce a new task to assess the model’s ability to propose molecules with underrepresented property values, assessing its capacity for out-of-distribution generalization.
-Our findings reveal shortcomings in existing structural metrics, which mistakenly validate open-shell molecules or molecules with invalid valence-charge configurations, underscoring the need for improved evaluation frameworks.
-Overall, this work paves the way for developing targeted property-guided generation methods, enhancing the design of molecular generative models for diverse applications.
-Keywords Molecule Generation  $\cdot$ Property Guidance  $\cdot$ Flow Matching  $\cdot$ SE(3) Equivariant Graph Neural Networks
+In this work, we introduce PropMolFlow, an approach for property-guided molecule generation based on geometry-complete SE(3)-equivariant flow matching.
+Integrating five different property embedding methods with a Gaussian expansion of scalar properties, PropMolFlow achieves competitive performance against previous SOTA diffusion models in conditional molecule generation while maintaining high structural stability and validity.
+Additionally, it enables faster sampling speed with fewer time steps compared to baseline models.
+We highlight the importance of validating the properties of generated molecules through DFT calculations.
+Furthermore, we introduce a task to assess the model’s ability to propose molecules with underrepresented property values, assessing its capacity for out-of-distribution generalization.
 Introduction
-Deep generative models show great promise to accelerate chemical discovery and drug design by direct statistical sampling of stable structures compared to resource-intensive physics-based approaches [1, 2].
-The past few years have witnessed rapid development of generative methods for molecule generation, such as variational autoencoders (VAEs) [3, 4], generative adversarial networks (GANs) [5], autoregressive models [6], score-based diffusion models [7, 8], flow-based models [9, 10, 11] and language models [12, 13].
-The predominant methods for molecule generation in 3D apply diffusion models combined with equivariant graph neural networks to generate molecules directly in 3D, providing a more accurate representation of molecular configurations compared to 1D SMILES and 2D graph-based representations [7, 14, 11, 8].
-Flow matching, in particular, has emerged as a flexible alternative to diffusion models, providing greater control over the choice of prior distributions and probability paths, as well as enhanced sampling efficiency [15, 16, 17].
-Its advantages have led to improved performance across various applications, including materials generation [18], protein design [19, 20], and molecule generation [21].
-Seminal examples of diffusion models and flow matching for molecule generation treated molecules as point clouds and adopted a continuous representation of the atom types, leveraging E(3)-equivariant graph neural networks (EGNN) to learn the score function in diffusion models and the conditional velocity field in flow matching [7, 21].
-However, the continuous representation of discrete molecular features, such as atom types and bond orders, fails to preserve their inherent discreteness, thus compromising their fidelity.
-Additionally, the EGNN architecture fails to capture molecular chirality, as an SE(3) generative process is required to account for chiral centers [22].
-Recent studies also highlight the importance of incorporating bond orders in molecular graphs to improve the stability and validity of generated molecules [23, 24].
-Building on these insights, recent diffusion models and flow matching approaches adopt a geometry-complete SE(3) generative process that incorporates bond orders into molecular graph representations, and utilize discrete diffusion or discrete flow matching for discrete molecular features such as atom types and bond orders [25, 26].
-Notably, FlowMol integrates these advances to achieve SOTA performance in unconditional molecule generation [26].
-However, its formulation remains limited to the unconditional setting, leaving its extension to property-guided molecule generation unaddressed.
-Conditional molecule generation involves designing new molecules that satisfy specific constraints or target properties.
-This is typically implemented by concatenating property values with node features to embed property information into molecular graphs, a technique that has been primarily demonstrated on the QM9 dataset [7, 8, 27, 25, 28].
-While this approach has proved effective, it may oversimplify the interaction between molecular graphs and properties.
-For example, Gebauer et al.
-suggest that applying a Gaussian expansion to a scalar property may help extract property information more robustly [29].
-Systematic studies of property embedding methods are lacking, and optimal embedding strategies may vary depending on the specific property.
-Moreover, previous methods use a pretrained classifier as the property predictor.
-Since this classifier is trained on the QM9 dataset, where all molecules are fully relaxed, it may struggle to make accurate predictions for generated molecules, which are likely not relaxed.
-Consequently, the reported results can be unreliable due to the limited generalization capacity of the pretrained classifiers/regressors used for evaluation.
-Last but not least, property-guided generation is typically evaluated in an ‘in-distribution’ setting, where property values sampled from within the QM9 distribution are used to generate molecules expected to exhibit the corresponding property values.
-In practice, for molecular discovery, we may want to evaluate the model’s ability to push the envelope of the original training distribution and generate molecules with underrepresented property values.
-In this work, we introduce PropMolFlow—a novel property-guided molecule generation framework—that integrates various property embedding methods with an SOTA SE(3)-equivariant flow matching process based on geometric vector perceptrons (GVP).
-We evaluate PropMolFlow on the QM9 dataset, showing that it outperforms previous diffusion models and flow matching methods, setting a new SOTA for property-guided molecule generation in 3D while achieving much faster inference times.
-We perform a systematic study of various property embedding methods and identify the optimal approach by property type and property evaluation tasks.
-We achieve SOTA performance of property-guided generation across many properties while maintaining SOTA performance for standard metrics, such as molecule stability and PoseBusters validity.
-We propose a new task for property-guided generation that evaluates PropMolFlow models’ capability in generating out-of-distribution molecules.
-In a first-of-its-kind analysis, we perform extensive density functional theory-based (DFT) calculations to validate the accuracy of our property-guided generation.
-We highlight the importance of DFT calculations for different tasks and properties, and release a curated DFT dataset of 10,773 molecules generated by PropMolFlow for broader community use.
-We propose “closed-shell ratio” and revised atomic and molecular stability as key metrics based on valence-electron configuration to filter out open-shell molecules and molecules with invalid valency–charge configurations, because the QM9 training data consists exclusively of closed-shell and charge-neutral molecules.
-Results
+Deep generative models are promising to accelerate chemical discovery by statistically sampling molecular structures, reducing reliance on costly physics-based simulations [sanchez-lengeling2018].
+State-of-the-art (SOTA) 3D molecule generation primarily uses diffusion with equivariant graph neural networks (EGNN), enabling accurate sampling of molecular geometries.
+[hoogeboom2022, satorras2022, xu2023].
+Flow matching is an emerging alternative to diffusion models [liu2022, lipman2023, albergo2023], offering flexibility over priors and probability paths while improving sampling efficiency across materials [hoellmer2025], proteins [campbell2024], and small molecules [dunn2024-2].
+Seminal diffusion [hoogeboom2022, xu2023] and flow-matching models [song2023] treat molecules as point clouds with continuous representations for discrete molecular modalities, and parameterize the generative process via E(3)-EGNN.
+However, continuous encodings disrespect inherently discrete features, such as atom types, and E(3) models cannot capture chirality [dumitrescu2024].
+Recent studies also found bond orders in molecular graphs to be important in improving validity of generated molecules [vignac2023, dunn2024-2].
+FlowMol [dunn2024-2] achieves SOTA unconditional molecule generation, by using geometry-complete SE(3) generative processes, explicit bond-order modeling, and discrete flow matching for atom types, formal charges and bond orders, though its extension to property-guided generation remains unaddressed.
+Property-guided molecule generation aims to design molecules that satisfy target properties, often by concatenating property values with node features in molecular graphs [hoogeboom2022, xu2023, morehead2024] — a strategy mainly explored on the single-molecule QM9 data [ramakrishnan2014quantum, wu2018moleculenet].
+Although effective, this approach may oversimplify how properties interact with molecular structures; for instance, Gebauer et al.
+[gebauer2022] suggest Gaussian expansions can extract property information more robustly.
+Systematic studies of property embedding methods are lacking, and optimal strategies may vary by properties.
+To validate properties of generated molecules, prior methods use separate property predictors trained on the fully relaxed QM9 molecules, which may struggle to make accurate predictions on non-relaxed, generated molecules.
+Moreover, property-guided generation is typically evaluated in an ‘in-distribution’ setting, where property values and atom counts sampled within the QM9 distribution are used to generate molecules [hoogeboom2022].
+Chemical discovery requires generating molecules with underrepresented or out-of-distribution (OOD) property values.
+We introduce PropMolFlow, a property-guided molecule generation framework that integrates various property embedding methods with an SE(3)-equivariant flow-matching process [dunn2024-2].
+We evaluate PropMolFlow on QM9, demonstrating its competitive performance versus previous methods, while achieving faster inference.
+Besides, we propose an out-of-distribution generation task, and validate properties of generated molecules using extensive density functional theory (DFT) calculations.
+Finally, we introduce the close-shell ratio and revised stability metrics to complement existing metrics and ensure more comprehensive evaluations for molecule generation.
+Results and Discussion
 Overview of PropMolFlow
@@
 A property embedding comprises an optional Gaussian expansion mapping (G.E.) followed by a multilayer perceptron (MLP) that projects a scalar property to a high-dimensional embedding space.
-Conditioning on the property is achieved by the interaction between the property embedding and node scalar features in the molecular graph.
-b, A joint flow matching process is used for molecule generation by learning different molecule modalities together.
-c, Five possible interaction types between a property embedding and a molecular graph.
-Node scalar features $[A,C]$ interact with a property embedding $P=\phi_{\mathrm{prop}}(k)$ via various operations.
-An MLP transformation $\varphi_{\theta}$ is applied to convert the dimension back to that of the original $[A,C]$, where necessary.
-$\odot$’ represents an element-wise Hadamard product, and $\oplus$’ stands for a ‘Concatenate’ operation.
+Conditioning on the property is achieved by the interaction between the property embedding and node scalar features.
+b, A joint flow matching process is used to generate different molecule modalities together.
+c, Five interaction types between a property embedding ‘P’ and a molecular graph.
+Node scalar features $A_{t},C_{t}]$ include atom type $A$ and formal charge $C$ at time $t$.
+An MLP transformation $\varphi_{\theta}$ is applied to convert the dimension back to that of the original $A_{t},C_{t}]$, where necessary.
+$\odot$’ represents an element-wise Hadamard product, and $\oplus$’ indicates a ‘Concatenate’ operation.
+$E_{t}$ and $X_{t}$ represent respective bond edge features and node position features.
 d, Gaussian expansion as augmented property embedding.
-Curves in the top panel correspond to five Gaussian basis functions that are evenly spaced between the minimum and maximum gap values.
-Centers of Gaussian basis functions are marked by gray dashed lines, and the red dashed lines represent the three example input values.
-The bottom panels show the function values for each Gaussian for two example inputs.
-In a molecule ball-and-stick representation, white, gray, red, and blue indicate H, C, O, and N, respectively.
-PropMolFlow builds on top of the state-of-the-art generative process using flow matching.
-A flow model learns the molecule graph distribution through training a velocity field conditioned on observed molecule samples.
-As shown in Figure 1a, PropMolFlow uses a fully-connected graph to represent a molecule and an SE(3)-equivariant graph neural network (GNN) to parameterize the velocity field.
-Property-guided generation is achieved through the interaction of a property embedding with the node scalar features of a sampled molecular graph.
-The property embedding is constructed by first transforming the raw property value using a fixed (non-trainable) Gaussian basis expansion, followed by a shallow multilayer perceptron (MLP) that produces the final trainable embedding.
-This Gaussian expansion is optional, and its utility depends on the specific property type and evaluation metric.
-Once trained, the SE(3)-equivariant graph neural network defines a joint flow matching process that denoises the molecular graph by simultaneously updating all modalities—such as atom types, bond orders, and 3D coordinates—to generate valid molecules (Figure 1b).
-Details of the joint flow matching are provided in Methods.
-We explored five specific operations, including ‘Concatenation’, ‘Sum’, ‘Multiply’, ‘Concatenate + Sum’, and ‘Concatenate + Multiply’, for combining property embeddings with node scalar features (Figure 1c).
-The dimension of node scalar features is preserved after interacting with the property embedding, providing a seamless integration into the original architecture.
-The non-trainable Gaussian expansion layer is used to enrich the property representations prior to being mapped to a property embedding via an MLP.
-Figure 1d illustrates an example of Gaussian expansion for the HOMO-LUMO gap.
-The top panel shows the five Gaussian basis functions and the bottom panel lists the function values for two example target properties, indicating the activation of each Gaussian basis function.
-We investigate the effectiveness of different property embedding strategies for six molecular properties in the QM9 dataset.
-The property guidance design space includes ten possible methods depending on whether the Gaussian expansion is used and which type of property embedding operation is employed.
-Details of the Gaussian expansion are provided in Methods.
-We trained PropMolFlow on the QM9 data with explicit hydrogen atoms [28, 30].
-The QM9 data contains 133,885 molecules, and each molecule contains up to 9 heavy atoms and consists of 3–29 atoms (average: 18 atoms), composed of up to 5 elements: C, N, O, F, and H.
-Molecules are fully optimized and molecular properties are obtained with DFT at the level of B3LYP/6-31G(2df,p).
-For property-guided generation, we considered six molecule properties: polarizability ($\alpha$), HOMO-LUMO gap, HOMO energy, LUMO energy, dipole moment ($\mu$), and heat capacity ($C_{v}$).
-To generate molecular graphs with explicit bond and charge information, we employed the QM9 SDF file distributed by DeepChem [30].
-Upon inspection, we discovered numerous inconsistencies: invalid bond orders and non-zero net charges in roughly 30,000 molecules, despite QM9’s requirement for charge-neutral, closed-shell valency [28].
-We corrected these discrepancies by reassigning bond orders or charges to enforce valency-charge consistency for the vast majority of entries and used the corrected data to train our PropMolFlow models.
-After the correction, we improved the structural validity of generated molecules significantly.
-Previous works that used bond orders overlooked this issue [23, 26]; instead, they modified evaluation metrics to tolerate chemically inconsistent bond–charge combinations, hence inflating molecule stability estimates contrary to QM9’s intended constraints.
-A detailed comparison between models using previous problematic data and the corrected data, together with the comparison between previous stability metrics and the revised ones, can be found in Supplemental Table LABEL:tab:data_and_metric_compare.
-A detailed description of our procedure and the alignment with the original QM9 XYZ data is provided in Supplemental Section LABEL:sec:qm9_data_fix.
-The revised QM9 SDF data is provided on ColabFit Exchange [31, 32].
-PropMolFlow achieves state-of-the-art property-guided generation
-We evaluated PropMolFlow based on its ability to accurately generate molecules with the target property, the validity of the generated molecules (i.e., whether they adhere to chemical rules), and the speed of generation.
-These assessments were first conducted under the in-distribution (ID) property-guided molecule generation task.
-The ID task involves sampling a range of property values from the same distribution as the training data.
-As such, it assesses the model’s capability to generate structures spanning a broad spectrum of target properties and has been widely adopted in prior studies of property-guided generative models [8, 21, 25, 27].
-In this task, a conditional generative model was first used to generate molecules conditioned on property values sampled from the same distribution as the training data (e.g., QM9).
-These sampled property values served as the input conditions.
-A separate property predictor/regressor was then applied to estimate the properties of the generated molecules.
-The predicted property values were compared to the original input conditions, and the mean absolute error (MAE) was used as the evaluation metric.
-A lower MAE indicates more accurate property conditioning and, therefore, better model performance.
-In PropMolFlow, we used an SE(3) GVP network for molecule generation and, correspondingly, trained a separate GVP regressor for each of the six molecular properties.
-For evaluation, we generated 10,000 molecules with input properties sampled from the property distribution of QM9 training data.
-Further details of the GVP regressors are provided in Methods.
-We compared PropMolFlow with five SOTA property-guided generative models, all trained on the QM9 dataset.
-Among these models, EEGSDE (Equivariant Energy Guided Stochastic Differential Equations) employs an energy function to guide the diffusion for property-guided generation [27]; EquiFM is the first equivariant flow matching model for 3D molecule generation [21]; GeoLDM introduces the first latent diffusion model for molecular geometry generation [8]; Geometry-Complete Diffusion Model (GCDM) constructs an SE(3) equivariant geometry-complete diffusion model using geometry-complete perceptrons for molecule generation [25, 12]; and JODO utilizes a diffusion graph transformer for the joint generation of 2D molecular graphs and 3D geometries.
-Overall, PropMolFlow achieves competitive performance against the SOTA models on property-guided generation (Table 1).
-It should be noticed that EEGSDE, EquiFM, GeoLDM, and GCDM do not include bond orders in their molecular graphs while JODO incorporates both bond existence and bond orders.
-All five baseline conditional models omit atomic charges, in contrast to PropMolFlow, which can generate molecules with non-zero atomic charges (see Methods for more details of the baseline models).
-The “Random (Upper-bound)” intuitive baseline fully shuffles property values, thereby removing any correlation between molecular structures and properties, and serves as an upper bound on error.
-The “# Atoms” baseline uses only the number of atoms as a predictor for molecular properties, capturing coarse size-dependent trends.
-The “QM9 (Lower-bound)” baseline corresponds to a pretrained EGNN regressor provided by Hoogeboom et al.
-[7], trained directly on QM9 and serving as a lower bound on achievable error.
-Improvement over the “Random” baseline indicates that the conditional generation effectively integrates property information.
-Surpassing the “# Atoms” baseline suggests that the generative model captures structural features beyond simple atom count when generating new molecules.
+Curves in the top panel correspond to five Gaussian basis functions that are evenly spaced between the minimum and maximum property values.
+Centers of Gaussians are marked by gray dashed lines, and the red solid lines represent two example inputs.
+The bottom panels show the function values for each Gaussian for two inputs.
+In a molecule configuration, white, gray, red, and blue indicate H, C, O, and N, respectively.
+PropMolFlow builds on top of the FlowMol architecture.
+[dunn2024-2] The model generates samples by integrating over NN-parameterized conditional velocity fields [lipman2023, albergo2023, liu2022].
+An SE(3)-EGNN based on geometric vector perceptrons (GVP) [jing2021] is used and molecules are represented by fully-connected graphs (Figure 1a).
+The generation is achieved by interacting property embeddings and node scalar features.
+Property embeddings are constructed by a non-trainable Gaussian expansion, followed with a shallow multilayer perceptron (MLP).
+This Gaussian expansion is optional, and its utility depends on property types and evaluation tasks.
+Once trained, the NN defines a joint flow matching that denoises the molecular graph by simultaneously updating all modalities—such as atom types, bond orders, and 3D coordinates—to generate molecules (Figure 1b).
+Details of flow matching and neural networks are provided in the ‘Joint Flow Matching’ and ‘Model Architecture’ sections.
+Five embedding methods and usage of Gaussian expansion are explored (Figure 1c,d).
+We evaluate the property-guided generation over six molecular properties in QM9.
+Details for property embeddings are provided in the ‘Property Embedding Operations’ and ‘Gaussian Expansion’ sections.
+PropMolFlow is trained on a revised QM9 SDF data with explicit hydrogen atoms, bond orders and formal charges [ramakrishnan2014quantum, wu2018moleculenet].
+This revised SDF file fixed the bond and charge inconsistencies of the original SDF file distributed by DeepChem [wu2018moleculenet].
+Data details are provided in the ‘Data’ section.
+Competitive property-guided generation with PropMolFlow
+We evaluated PropMolFlow on its ability to generate molecules with target properties, molecular structural validity, and inference speeds.
+These assessments were conducted under the in-distribution (ID) task.
+This task assesses the model’s capability to generate structures spanning a spectrum of joint target properties and atom counts within data distributions and has been adopted in prior studies [xu2023, song2023, morehead2024, bao2023].
+Generative models sample molecules by property inputs, and a separate property predictor evaluates properties of generated molecules.
+Predicted values are compared to input values, and the mean absolute error (MAE) was used as the evaluation metric.
+A lower MAE indicates better model performance.
+Further details of how to sample molecules and GVP regressors are provided in the ‘Inference’ and ‘GVP regressor details’ sections.
+We compared PropMolFlow with five baseline models, including EEGSDE [bao2023], EquiFM [song2023], GeoLDM [xu2023], GCDM [morehead2024] and JODO [huang2024], all trained on QM9.
+Two intuitive baselines are also included to represent the respective upper and lower bounds of MAEs, namely “Random (Upper-bound)” and “QM9 (Lower-bound)”, along with another baseline “# Atoms” (see the ‘Baselines’ section for more details).
 Table 1: Performance of PropMolFlow with respect to property alignment.
-Results for the baseline models use an equivariant graph neural network (EGNN) as the regressor, whereas PropMolFlow results use our pretrained SE(3) equivariant GVP regressor.
-Best results are bolded, whereas the second-best results are underlined.
+Results for the baseline models use an equivariant graph neural network (EGNN) as the regressor, whereas PropMolFlow results use our pretrained SE(3) GVP regressor as well as an EGNN regressor.
 JODO results were reported on our sampled molecules using publicly available model checkpoints.
 Property $\alpha$ $\Delta\epsilon$ $\epsilon_{\rm{HOMO}}$ $\epsilon_{\rm{LUMO}}$ $\mu$ $C_{v}$
-Units Bohr3 meV meV meV Debye cal/(mol $\cdot$ K)
+[…]
 QM9 (Lower-bound) 0.10 64 39 36 0.043 0.040
@@
 JODO 1.44 333 231 260 0.620 0.580
-PropMolFlow 1.31 391 254 315 0.620 0.626
-Notably, PropMolFlow shows the lowest MAE for two properties—$\alpha$ and $\mu$—and comparable performance for $C_{v}$ and $\epsilon_{\rm{HOMO}}$ versus the SOTA JODO, and the second best performance for the other two properties $\epsilon_{\rm{LUMO}}$ and $\Delta\epsilon$.
-For instance, PropMolFlow achieves substantially lower MAEs compared to GCDM for $\epsilon_{\rm{LUMO}}$ (315 meV vs.
-479 meV), and $\mu$ (0.620 Debye vs.
-0.844 Debye), demonstrating a clear advantage in controllable molecular generation over prior diffusion- and flow-based approaches.
-The superior performance of PropMolFlow can be ascribed to the inclusion of bond order information and discrete treatment for atom types, charges and bond orders, which encodes structural information more accurately and hence offers a more robust structure–property mapping for generated molecules.
-It also benefits from the thorough exploration of property-dependent optimal embedding methods (Supplementary Table LABEL:tab:mae_range and Supplemental Section LABEL:sec:model_stats).
-For instance, the MAE variation for $\Delta\epsilon$ can vary by more than 40%, from 391 meV using a ‘Concatenate_Sum’ embedding to 551 meV using a ‘Concatenate’ embedding, both without Gaussian expansions.
+PropMolFlow, GVP 1.31 391 254 315 0.620 0.626
+PropMolFlow, EGNN 1.36 391 246 312 0.632 0.640
+Overall, PropMolFlow achieves competitive performance against the SOTA models (Table 1).
+It shows the lowest MAE for $\alpha$ and $\mu$, and comparable performance for $C_{v}$ and $\epsilon_{\rm{HOMO}}$ versus the SOTA JODO, and the second best performance for $\epsilon_{\rm{LUMO}}$ and $\Delta\epsilon$.
+The superior performance of PropMolFlow benefits from the thorough exploration of property-dependent optimal embedding methods (Supplementary Table LABEL:tab:mae_range and Supplementary Section LABEL:sec:model_stats).
+For instance, the MAE for $\Delta\epsilon$ can vary by more than 40%, from 391 meV using a ‘Concatenate_Sum’ embedding to 551 meV using a ‘Concatenate’ embedding, both without Gaussian expansions.
+To avoid bias of comparison using different predictors, we also report MAEs using EGNN predictors released by Bao et al.
+[bao2023] in Table 1.
+It confirms that different predictors produce consistent MAEs.
 Rapid molecule generation with high structural fidelity
-Figure 2: Chemical validity and sampling efficiency of PropMolFlow against five baseline models.
-a, Molecule stability.
-b, RDKit validity.
-c, Uniqueness.
-d, PoseBusters validity.
-e, Closed-shell ratio.
-f, Sampling time for 10k molecules.
-Mean $\pm$ standard deviations are reported across six molecule properties.
-Structural validity metrics are shown in the value range between 0.4 and 1, while sampling times are in the unit of minutes.
-Besides the ability to generate molecules with target properties, the structural validity and speed of generation are also critical in ensuring the success and robustness of chemical discovery with generative models.
-We further evaluated the performance of PropMolFlow by assessing the structural validity of the generated molecules and the sampling efficiency.
-Structural validity refers to whether the generated molecules conform to basic chemical rules.
-This was measured using several metrics, including atomic stability, molecule stability, RDKit validity, ‘uniqueness and validity’, PoseBusters validity, and closed-shell ratio.
-Atom stability is given by the proportion of atoms with correct valency.
-To account for atomic charges, an atom is considered to have the correct valency if its formal charge balances its explicit valency, as defined in RDKit [33].
-For instance, a nitrogen atom should carry a +1 formal charge if it has a valency of 4.
-Molecule stability defines the proportion of molecules in which all atoms are stable and the molecule is charge neutral (zero net formal charges).
-RDKit validity refers to the ratios of molecules that pass RDKit’s sanitization check.
-‘Uniqueness and validity’ is defined as the proportion of molecules that are RDKit valid and unique in their SMILES representation.
-PoseBusters validity corresponds to the fraction of molecules that pass many de novo chemical and structural validation tests, such as all atom connectivity, valid bond lengths and angles, and absence of internal clashes [34].
-A molecule is closed-shell valid if it has an even number of valence electrons.
-Since all molecules in QM9 are closed-shell, any generated open-shell species are considered invalid.
-Sampling efficiency is given by the wall-clock time required to generate 10,000 molecules.
-Computational settings for sampling are provided in Methods.
-Across nearly all stability and validity benchmarks, PropMolFlow consistently outperforms baseline modes (Figure 2a–e).
-The only exception is the closed-shell ratio, where it slightly underperforms JODO—a diffusion model that omits atomic formal charges and adds additional bond features such as aromaticity existence.
-Moreover, PropMolFlow achieves high chemical validity with much faster sampling efficiency.
-Flow matching is normally much faster than diffusion models due to its shorter, deterministic probability path (Eq.
-(6) and (8)) and optimal transport (Supplementary Section LABEL:sec:ot[35, 21].
-PropMolFlow requires only 100 time steps compared to the 1000 diffusion time steps, and achieves a speedup for sampling of at least 8$\times$ over diffusion-based models and it is nearly 2$\times$ faster than EquiFM (Figure 2f).
-When evaluating metrics on a per-property basis, PropMolFlow consistently achieves the highest molecular stability and validity across all properties (Supplementary Tables LABEL:tab:propmolflow_metric_by_property-LABEL:tab:jodo_metric_by_property).
-Across properties, PropMolFlow’s metrics follow the order: $\mu\approx\Delta\epsilon>\epsilon_{\mathrm{HOMO}}>\epsilon_{\mathrm{LUMO}}>\alpha>C_{v}$.
-This trend generally holds for the baseline models as well, with the exception of EEGSDE, suggesting a more complex structure–property relationship for $\alpha$ and $C_{v}$.
-This complexity may be due to their strong dependence on molecular geometry and the limitation of the training data, which includes only fully relaxed molecules.
-Across metrics, molecular stability and “Validity & Uniqueness” exhibit over 2% variations across properties, whereas molecular validity and PoseBusters validity vary by less than 1%.
-Besides, all methods in this study generate a non-negligible fraction of open-shell molecules: on average 1.6% for JODO and at least 4% for the other methods.
-This highlights a previously overlooked issue in generative molecular modeling: the inability to consistently respect the valence electron constraints inherent to the training data.
-Addressing this limitation is essential for improving the chemical validity of generated molecules especially when atomic formal charges are considered in molecular graphs, and future work should improve upon this direction—perhaps by explicitly including closed-shell constraints, as we do here by adding bond orders into molecular graphs—to ensure better alignment with the reference distribution.
-Property predictor as DFT proxy reveals systematic inductive bias
-Figure 3: Performance of GVP property predictors for PropMolFlow-generated molecules without and with relaxation.
-a, Comparison between Target, DFT and GVP shows the reliability of GVP to evaluate the MAE metrics commonly used for property-guided generation.
+Structural validity and inference speeds are critical in success of chemical discovery with generative models.
+Structural validity refers to whether generated molecules conform to chemical rules.
+This was measured using five metrics, including atomic stability, molecule stability, RDKit validity [rdkit], ‘uniqueness and validity’, PoseBusters validity [buttenschoen2024], and closed-shell ratio.
+Inference speeds are evaluated by the wall-clock time for generating 10,000 molecules.
+Details of structural metrics and computational settings for sampling are provided in respective ‘Evaluation metrics’ and ‘Computational settings’ section.
+[…]
+[…]
+Systematic inductive bias of property predictors
+Figure 3: Performance of GVP property predictors without and with DFT relaxation.
+a, Comparison between Target, DFT and GVP shows the reliability of GVP in evaluating MAE metrics commonly used for property-guided generation.
 Comparison between GVP and GVP-R, and between DFT and DFT-R shows the structural dependence of GVP-predicted and DFT-predicted properties, respectively.
-Comparison between GVP and DFT with and without relaxation shows the reliability of GVP in capturing the ground-truth DFT values over raw and relaxed structures.
-Target is the input target property values used by PropMolFlow models to generate models.
-‘DFT’ and ‘GVP’ are respective DFT-computed and GVP-evaluated property values on directly generated raw molecules.
+Comparison between GVP and DFT with and without relaxation shows the reliability of GVP in capturing ground-truth DFT values for both raw and relaxed structures.
+‘Target’, ‘DFT’ and ‘GVP’ denote input, DFT-calculated and GVP-predicted property values on raw molecules, respectively.
 ‘DFT-R’ and ‘GVP-R’ refer to values evaluated on DFT-relaxed molecules.
+The $d$ indicates the MAE distances between two property-value vectors.
 b, Pairwise comparison between Target, DFT and GVP on raw molecules.
 c, GVP versus DFT for both raw and DFT-relaxed molecules.
-d, Structure variation and property dependence on the DFT relaxation.
-Molecules with the highest RMSD for each property and their DFT and GVP values are listed as examples.
-In molecule representation, gray, red, blue, and white colors indicate C, O, N, and H, respectively.
-Property values for $\alpha$, $\Delta\epsilon$, and $C_{v}$ are in units of Bohr3, eV, and cal/(mol $\cdot$ K), respectively.
-A separate surrogate model is needed to evaluate the properties of generated molecules.
-Typically, this surrogate model shares the same architecture as the generative neural network model [7].
-However, using models from the same family may introduce inductive biases, making it necessary to validate results using external property predictors, such as DFT.
-To this end, we performed DFT calculations on 1000 molecules selected from the 10,000 structures generated for each property in the ID task.
-For directly generated molecules, multi-step filtering is essential to ensure that the structures are chemically valid and suitable for downstream applications.
-We filtered the 1000 molecules by five criteria: Molecule stability, RDKit validity, PoseBusters validity, closed-shell validity, and multi-fragment check.
-The multi-fragment check filters out molecules that with more than one disconnected fragment.
-For molecules passing all filters, property values are evaluated using both a GVP property predictor and DFT calculations.
-We also filtered out molecules whose DFT calculations did not converge, resulting in 90% of the molecules remaining for the subsequent analyses (Table LABEL:tab:no_sp_dft).
-To evaluate the reliability of GVP property predictor and the impact of structural relaxation, we conducted a comprehensive comparison among target property values (Target), DFT-calculated properties (DFT), and GVP predictions (GVP)—using both directly generated and DFT-relaxed molecules (Figure 3a).
-We chose three example properties, $\alpha$, $\Delta\epsilon$ and $C_{v}$, for the following analyses because the results for the other three properties are similar to those of $\Delta\epsilon$ (Supplemental Figure LABEL:fig:si_target_dft_gvp).
-Example orders of pairwise MAE distances for $\alpha$ are also provided in Figure 3a.
-We first assess the reliability of GVP on the initially generated molecules (Figure 3b).
-The MAEs between the target values and DFT (Target vs.
-DFT) are comparable to those betwen the target and GVP predictions (Target vs.
-GVP) for $\alpha$ and $\Delta\epsilon$, suggesting that GVP property predictors are statistically reliable for estimating properties of directly generated molecules.
-However, GVP consistently underestimates the MAEs, revealing an inductive bias in its predictions.
-Notably, although the MAEs of GVP and DFT relative to the target are similar, the MAEs between GVP and DFT themselves are significantly larger—for example, 0.91 vs.
-0.1 Bohr3 for $\alpha$, and 204 vs.
-43 meV for $\Delta\epsilon$.
-For $C_{v}$, the MAE between the target values and DFT is much larger than that between the target values and GVP.
-However, this discrepancy is not due to poor reliability of the GVP property predictors, but rather stems from an issue with DFT-calculated heat capacities ($C_{v}$) on unrelaxed structures.
-DFT-computed $C_{v}$ values depend on vibrational frequencies, which are highly sensitive to molecular geometry.
-To validate this analysis, we relaxed the structures, resulting in the DFT vs.
-Target MAE dropping from 1.68 to 0.68 cal/(mol $\cdot$ K), closely aligning with the MAE of 0.61 cal/(mol $\cdot$ K), predicted by GVP.
-Evaluation on directly generated molecules without geometry optimization reflects the statistical quality of the learned structure–property relationships captured by both PropMolFlow models and the GVP predictors.
-However, from a chemical discovery perspective, it is essential to relax the structures and evaluate properties based on the relaxed chemical stable conformations that are more likely to occur in real-world experiments.
-To assess the reliability of GVP on relaxed structures, we compare its predictions with DFT-computed property values for both unrelaxed and relaxed molecules (Figure 3c).
-For $C_{v}$, a much closer agreement between DFT and GVP was observed on relaxed structures.
-In contrast, for $\alpha$, a notable offset between GVP and DFT appears when using relaxed structures, indicating that GVP is more reliable on unrelaxed structures.
-This can be attributed to the inductive bias of GVP property predictors—specifically, GVP tends to perform better on directly generated structures by PropMolFlow, as both models share the same architecture.
-This bias is further supported by the increased MAEs between GVP and DFT after relaxation for all properties except $C_{v}$ (Supplemental Figure LABEL:fig:id_mae_heatmap).
-For $\Delta\epsilon$, despite a slight broadening of the peak at lower $\Delta\epsilon$ values is observed, both DFT and GVP predictions remain largely unchanged, suggesting that this property has weak dependence on structural relaxation.
-To further quantify the effect of relaxation, we calculated the symmetry-aware root mean square distances (RMSDs) between atomic positions of raw structures and DFT-relaxed structures.
-Given the same chemical compositions and element-type ordering, the RMSD between any two molecules (M0 and M1) is the minimum distance between these two structures considering the translational and rotational symmetry operations, which can be expressed as:
-$\mathrm{RMSD}(\rm{M}_{0},\rm{M}_{1})=\min_{P\in SE(3)}\{d\left(M_{0},P(M_{1})\right)\}$ (1)
-Where $SE(3)$ represents the group of operations that respect the translational and rotational symmetry.
-The RMSD distributions for $\alpha$, $\Delta\epsilon$ and $C_{v}$ are showed in Figure 3d.
-All RMSD distributions are highly left-skewed and concentrated around 0 Å, meaning the majority of PropMolFlow-generated structures are close to their relaxed, stable counterparts.
-To estimate the property sensitivity to structural relaxation, we selected the structures whose RMSDs are larger than 0.03 Å—RMSDs lower than this threshold are considered very close to the relaxed structures, and in Eq.
-(2), we define the normalized structure sensitivity of DFT-calculated properties to allow comparisons between different properties.
-$\chi_{\rm{DFT}}=\frac{|{\delta q_{\rm{DFT}}|}}{(q_{0.99}-q_{0.01})\cdot\rm{RMSD}}$ (2)
-where $\delta q_{\rm{DFT}}$ is the change of DFT-calculated property upon relaxation, and the normalizer ($q_{0.99}-q_{0.01}$) is the difference between the 99th quantile and the 1th quantile of the property distribution of QM9 training data.
-This range is chosen over the min-max range because it is robust to outliers.
-Despite similar RMSD distributions across the three properties, the differences are more pronounced in the structural dependence of DFT-calculated properties, following the order of $C_{v}>\alpha>\Delta\epsilon$.
-On the right panel of Figure 3d, molecule configurations with the highest RMSDs are shown, and their RMSD, taget input, DFT-calculated, GVP-predicted property values are listed to rationalize the relationship between GVP, DFT and Target values with and without relaxation.
-DFT property values for relaxed structures can be either closer or farther to the target input values.
-For these molecules with large structural changes, the GVP predictions are nearly unaltered, showing a much weaker structural dependence compared to that of DFT.
-Although DFT property changes are more significant than GVP upon relaxation for the three extreme cases, GVP predictions are close to DFT-calculated values for both raw and relaxed molecules, offering reliable evaluation of PropMolFlow models and robust guidance towards chemical discovery.
-Interpolated structures follow chemically intuitive trends
-To assess whether PropMolFlow models have learned a continuous and smooth structure–property relationship, we performed interpolation by varying the target property values while fixing the number of atoms in the generated molecules.
-For each property value, ten molecules were sampled, and the one that was chemically valid and whose DFT-calculated property most closely matched the target property value was selected.
-For each property, we show the minimum and maximum target values along with their DFT-calculated values below the molecular configurations.
-The target property value, DFT-calculated value, and GVP predictions for each configuration are provided in Supplemental Table LABEL:tab:interpolation.
-Figure 4: PropMolFlow generated molecules when interpolating among different property values while fixing the number of atoms at 19 for sampling.
-The minimum and maximum target properties in red and corresponding DFT-calculated properties in blue are provided below the configurations.
-All molecules chosen here pass the five filtering criteria and have DFT values closest to the target properties.
-C, H, O, and N are in gray, white, red, and blue colors, respectively.
-Property units are provided in the square brackets under each property symbol.
+d, Root mean squared distances (RMSDs) and normalized property sensitivity due to DFT relaxation.
+Molecules with the highest RMSDs for each property and their corresponding DFT and GVP values are shown.
+In molecular representations, gray, red, blue and white indicate C, O, N and H atoms, respectively.
+Property values for $\alpha$, $\Delta\epsilon$ and $C_{v}$ are in units of Bohr3, eV and cal mol-1 K-1, respectively.
+A property predictor is needed to evaluate the generated molecules.
+Typically, this predictor shares the same architecture as the generative model, which may introduce inductive biases in predictions.
+In this work, DFT calculations were carried out on selected and filtered molecules and compared to input (‘Target’) and GVP-predicted values, without and with (‘-R’) structural relaxations.
+For example, we show orders of pairwise MAE distances for $\alpha$ in Fig.
+3a.
+We show results for $\alpha$, $\Delta\epsilon$ and $C_{v}$ in Fig.
+3b–d, and for the other properties in Supplementary Figure LABEL:fig:si_target_dft_gvp, which are similar to $\Delta\epsilon$.
+DFT settings are included in the ‘DFT’ section, the filtering procedure are provided in the ‘Molecule filtering’ section, and numbers of filtered molecules are listed in Supplementary Table LABEL:tab:no_sp_dft.
+First, we examine GVP on the raw generated molecules (Fig.
+3b).
+MAEs of “Target vs.
+DFT” are comparable to “Target vs.
+GVP” for $\alpha$ and $\Delta\epsilon$, suggesting that GVP predictors are reliable for estimating “Target vs.
+DFT” MAEs.
+However, GVP consistently underestimates the MAEs, revealing an inductive bias.
+For $C_{v}$, the “Target vs.
+DFT” MAE is much larger than “Target vs.
+GVP”.
+This is because DFT-computed $C_{v}$ depends on vibrational frequencies, which are highly sensitive to geometry.
+Relaxation resolves this issue, reducing the MAE to 0.68 cal/(mol $\cdot$ K), still higher than, but closer to “Target vs.
+GVP” (0.61).
+For chemical discovery, it is also important to relax the structures.
+Fig.
+3c shows a much closer agreement between DFT and GVP after relaxation for $C_{v}$, while a notable offset is observed for $\alpha$.
+It indicates that GVP is more reliable on unrelaxed structures, perhaps due to its inductive bias.
+This is also true for other properties (Supplementary Figure LABEL:fig:id_mae_heatmap).
+To further quantify the relaxation effect, we calculated the root mean square distances (RMSDs) between atomic positions before and after relaxation.
+Fig.
+3d shows that RMSD distributions are left-skewed toward 0 Å, meaning most unrelaxed structures are close to their relaxed counterparts.
+To estimate the property sensitivity to relaxation, we calculated the normalized property sensitivity to allow comparisons between different properties.
+Definitions of RMSDs and normalized property sensitivity are provided in the ‘Evaluation metrics’ section.
+Despite similar RMSD distributions, the differences are more pronounced in the property sensitivity, following the order of $C_{v}>\alpha>\Delta\epsilon$.
+Fig.
+3d also shows configurations with the highest RMSDs, their target, DFT-calculated, and GVP-predicted values.
+DFT values with relaxation can be either closer or farther to Target.
+Despite high RMSDs, GVP values are nearly unaltered with relaxation, showing a much weaker structural dependence of GVP versus DFT.
+Overall, GVP predictions are close to DFT values for all cases, hence it can offer a statistically reliable evaluation of PropMolFlow models despite the inductive biases.
+Interpolated structures aligned with chemical intuition
+To assess whether PropMolFlow has learned a smooth structure–property relationship, we performed interpolation by varying target values while fixing atom counts of molecules at 19.
+For each value, ten molecules were sampled, and the one whose DFT-calculated property is closest to the input was selected.
+We show the configurations for selected molecules, and minimum and maximum target values of each property along with corresponding DFT values in Fig.
+4.
+Numeric results for all configurations are provided in Supplementary Table LABEL:tab:interpolation.
+Figure 4: Interpolation study by varying property values.
+The minimum and maximum target properties (red) and the corresponding DFT-calculated properties (blue) are shown below each configuration.
+All molecules shown pass the filtering criteria and have DFT values closest to the target properties.
+In molecular representations, gray, white, red and blue indicate C, H, O and N atoms, respectively.
+Property units are provided in square brackets under each property symbol.
 […]
-The interpolated structures generated by PropMolFlow exhibit smooth and chemically intuitive transitions as property values vary, demonstrating that the model has learned meaningful structure—property relationships.
-These results highlight the model’s ability to generate property-guided molecules that align with established chemical principles.
-Figure 5: Performance of PropMolFlow in OOD generation.
-a, Distribution of DFT calculated and GVP predicted values for PropMolFlow generated molecules, and the property distribution of QM9 is also included.
-The vertical black dashed line in histograms represents the target property value $q_{0.99}$.
-Curves on top of histograms are fitted with a kernel density estimation.
-b, Three example molecules that do not exist in QM9 but are found in a larger PubChem dataset are included in the left panel.
-Numbers below the configurations are DFT calculated property values on raw molecules generated by PropMolFlow models.
-C, H, O, N, and F are in gray, white, red, blue, and yellow colors, respectively.
-Property values for $\alpha$, $\Delta\epsilon$, and $C_{v}$ are in units of Bohr3, eV, and cal/(mol $\cdot$ K), respectively.
-PropMolFlow pushes molecule generation toward out-of-distribution property space
-Although ID task is typically used to evaluate property-guided generative models, it is also of practical interest to test molecular generation for properties that are underrepresented in the training dataset.
-To this end, we propose an ‘out-of-distribution’ (OOD) task designed to assess the model’s ability to generate molecules conditioned on property values that are rare or sparsely represented in the original training distribution.
-In this task, we chose the 99th quantile of the property distribution of QM9 training data as the target property value.
-The target value for each property is provided in Supplemental Table LABEL:tab:prop_choice.
-We sampled 1,000 structures conditioned on the chosen property value using the models with top-performing embedding methods (Supplemental Table LABEL:tab:best_model).
-These generated molecules then pass through the same multi-step filtering procedure used in the ID task, and their properties are evaluated by both GVP and DFT.
-Here we compare the property distribution of generated molecules against that of the QM9 training data for $\alpha$, $\Delta\epsilon$ and $C_{v}$ (Figure 5a).
-Results for other properties are provided in Supplemental Figure LABEL:fig:si_ood.
-The DFT results demonstrate that the OOD generation successfully shifts the property distribution of generated molecules towards the specified target values (indicated by the dashed line), and the DFT-predicted distribution is nearly perfectly centered on the target property value for the polarizability ($\alpha$).
-However, for $\Delta\epsilon$ and $C_{v}$, the DFT-predicted distribution is off the target by a small margin.
-For example, both the GVP and DFT distributions for the gap are centered around 8.95 eV, which is 0.41 eV away from the target value of 9.36 eV.
-This deviation is likely due to the scarcity of training data in the corresponding region of property space.
-One should note the DFT issue for $C_{v}$ on unrelaxed structures which can be largely addressed by performing relaxation.
-While property guidance effectively steers the generative model toward the desired values, the model’s capacity to extrapolate is ultimately constrained by the coverage of the training distribution.
-Other guidance approaches, such as classifier-free guidance [36], can adjust the weights for the property conditioning, potentially achieving better alignment with target property values, as demonstrated in the conditional generation of inorganic crystal materials [37].
-Moreover, the GVP distributions are almost overlapped with the DFT ones, showing desired extrapolation capability of the GVP property predictors.
-To further evaluate the novelty of generated molecules, Figure 5b presents three example molecules that are not part of the QM9 dataset but are found in the larger PubChem database [38].
-These molecules exhibit the DFT properties close to their target input values.
-This demonstrates PropMolFlow’s ability to generate chemically valid molecules with desired target properties that lie outside the original training domain.
-Conclusion
-In this work, we introduce PropMolFlow, a property-guided molecule generation framework based on the SOTA SE(3) flow matching.
-Our approach outperforms existing diffusion and flow matching models in the conditional generation setting, offering improved property alignment, structural validity and sampling efficiency across many molecular properties.
-We conduct a systematic evaluation of property embedding strategies and identify optimal configurations that depend on both the target property and the task regime—whether ID or OOD.
-To enable a more comprehensive evaluation, we propose two new metrics and one new task: The closed-shell ratio and revised atomic stability and molecular stability to account for charged species, which quantify how well the generated molecules align with the training data in terms of valency–charge configurations, and an OOD task, designed to evaluate controllability in sparsely represented regions of property space.
-To validate the fidelity of generated molecules and reliability of property predictors, we perform extensive DFT calculations.
-PropMolFlow generates molecules that are close to the stable, relaxed counterparts.
-Generally, GVP property predictors provide reliable approximations of the true DFT MAEs against target values despite the relative large difference between GVP and DFT.
-We demonstrated the inductive bias of property predictors, which lead to a systematic underestimation of true MAEs and its partial failure when predicting on unseen relaxed structures.
-Additionally, we curated a DFT dataset of generated molecules, which is made available through the ColabFit Exchange [31, 39].
-While this work focuses on small molecules from the QM9 dataset, PropMolFlow is readily extendable to larger molecular datasets with more atoms, as well as to materials datasets, as long as property labels are available.
+Toward out-of-distribution generation
+We propose an OOD task to test generation for underrepresented property values.
+The 99th quantile of the training data distribution was chosen as the target value and the atom counts were chosen according to the ‘Inference’ section.
+Specific target values are provided in Supplementary Table LABEL:tab:si_prop_val_stats.
+1,000 molecules were sampled with top-performing models in this task (Supplementary Table LABEL:tab:best_model), and we filtered structures using the same procedure as that used in the ID task.
+We compare DFT and GVP property distributions against that of the QM9 training data (Fig.
+5a).
+DFT distributions shift toward the targets (dashed vertical lines).
+The distribution mode is very close to the target for $\alpha$.
+For $\Delta\epsilon$ and $C_{v}$, DFT distribution modes are off the targets by a small margin, for example 0.41 eV for $\Delta\epsilon$ (8.95 versus target 9.36 eV).
+This deviation is likely due to the scarcity of training data in the corresponding property region; when conditioning on the 50th quantile (medians), modes are closer to the targets (Supplementary Figure LABEL:fig:si_q05_vs_q099).
+Moreover, GVP distributions are almost overlapped with DFT ones, suggesting their desired extrapolation performance.
+To examine the novelty of generated molecules, Fig.
+5b presents three generated molecules whose SMILES representations are not part of QM9 but are found in the larger PubChem database [kim2021].
+We also evaluated the novelty of generated molecules by Tanimoto similarity to the training data.
+Fig.
+5c indicates that 20–35% of the molecules exist in the training data.
+Using 0.8 as a cutoff, novelty ratios are 58–75% across all properties, demonstrating substantial exploration beyond the training data distribution.
+Results for other properties are provided in Supplementary Figure LABEL:fig:si_ood, and additional extrapolation to more extreme property values or atom counts can be found in respective Supplementary Figure LABEL:fig:si_various_values and Supplementary Figure LABEL:fig:si_ood_no_atoms.
+Figure 5: Toward out-of-distribution generation.
+a, Distribution of DFT-calculated and GVP-predicted property values for molecules generated by PropMolFlow; the property distribution of the QM9 training data is also shown.
+The vertical black dashed line in the histograms denotes the target property value $q_{0.99}$, corresponding to the 99th percentile of the training-data distribution.
+Curves overlaid on the histograms are kernel density estimation fits.
+b, Three example molecules absent from QM9 but present in a larger PubChem dataset are shown in the left panel.
+Numbers below the configurations indicate DFT-calculated property values for raw molecules generated by PropMolFlow.
+In molecular representations, gray, white, red, blue and yellow indicate C, H, O, N and F atoms, respectively.
+Property values for $\alpha$, $\Delta\epsilon$ and $C_{v}$ are in units of Bohr3, eV and cal mol-1 K-1, respectively.
+c, Maximum Tanimoto similarity between generated, filtered molecules and the training data computed using Morgan fingerprints.
+Dashed lines indicate the 0.8 similarity cutoff used to define novel molecules.
+Discussion
+While PropMolFlow can generate chemically valid and novel molecules with strong property alignment and fast inference, its ability to extrapolate could be further enhanced, potentially by integrating the model into an active-learning or reinforcement-learning framework.
+The current implementation is limited to the small-molecule QM9 dataset; extending PropMolFlow to larger molecular datasets and conditioning on properties of greater practical relevance represents an important future direction.
+In some cases, the property distribution of generated molecules exhibits deviations from target values.
+Property alignment could be improved by incorporating advanced guidance mechanisms with tunable guidance weights [ho2022, karras2024].
+Moreover, despite faster inference than diffusion models, flow matching still requires multi-step integration .
+Additional efficiency gains may be possible by learning flow maps that support direct jumps along probability paths [boffi2025].
+However, these approaches have shown effectiveness only on continuous data; learning flow maps for multi-modal 3D molecules remains a challenge.
+Although extending PropMolFlow to multiple properties is conceptually straightforward, the optimal strategy for synthesizing different properties and for managing potential correlations remains unclear.
+Another challenge concerns conformational stability: PropMolFlow does not guarantee the generation of the most stable conformers.
+Injecting energies into de novo generation is an open question, though recent advances in conformer generation with fixed compositions, such as adjoint Schrödinger bridge sampling [liu2025], may provide useful insights.
+Furthermore, several techniques, such as fake atoms and self-conditioning, were recently introduced in the unconditional FlowMol model [dunn2025-3] while this work was under development, and may further benefit the PropMolFlow conditional model.
+Despite these limitations and challenges, PropMolFlow demonstrates reliable generative modeling in chemical discovery and provides a strong foundation for future advancements.
 Methods
-Joint Flow Matching for Molecule Generation
-Ideally, new molecules would be generated by sampling directly from the probability distribution over valid molecular graphs $q(x)$.
+Joint flow matching
+Molecules are represented by fully-connected graphs $G$.
+Graph nodes encode the atomic types $A$, charges $C$ and positions $X$ of a molecule, and graph edges include bond orders $E$.
+Ideally, new molecules would be generated by sampling directly from the probability distribution over valid molecular graphs $q(x)$, where $x$ is an arbitrary sample.
 However, this target distribution lacks a tractable analytical form and can only be approximated through a generative process.
 To this end, we adopt flow matching to learn an approximation of $q(x)$.
-This flow matching is parameterized by an SE(3) equivariant graph neural network built on top of FlowMol architecture [26], details of which are provided in Supplemental Section LABEL:sec:appendix_model.
-In flow matching, the objective is to learn a time-dependent velocity field $u_{t}^{\theta}$ parameterized by a neural network $\theta$, which defines a probability path $p_{t}(x):[0,1]\times x_{0}\rightarrow x_{1}$ from a base distribution, $p_{0}(x)$ to a target distribution, $p_{1}(x)=q(x)$.iiiWe use subscripts to denote time dependence, e.g., $p_{t}(x)$ The flow induced by the velocity field is given by the ODE $\dot{x}=u_{t}\left(x\right)$, which describes the evolution of samples along the probability paths.
+This flow matching is parameterized by an SE(3) equivariant graph neural network built on top of FlowMol architecture [dunn2024-2], details of which are provided in the subsequent ‘Model architecture’ section.
+In flow matching, the objective is to learn a time-dependent velocity field $u_{t}^{\theta}$ parameterized by a neural network $\theta$, which defines a probability path $p_{t}(x):[0,1]\times x_{0}\rightarrow x_{1}$ from a base distribution, $p_{0}(x)$ to a target distribution, $p_{1}(x)=q(x)$ (subscripts denote time dependence, for example, $p_{t}(x)$) The flow induced by the velocity field is given by the ODE $\dot{x}=u_{t}\left(x\right)$, which describes the evolution of samples along the probability paths.
 Designing a flow matching process thus involves two key steps: 1) specifying a probability path that satisfies the boundary conditions $x_{0}\sim p_{0}$ and $x_{1}\sim p_{1}$, and 2) training a neural network $\theta$ to approximate the velocity field that transports samples along this path.
 In practice, while accurately modeling the exact velocity field is challenging, conditioning it on a random variable $z$ leads to a more tractable formulation.
-Importantly, optimizing the conditional velocity field is equivalent to optimizing the marginal velocity field [40].
-The endpoint reformulation of [24] reparametrizes the flow matching problem by focusing on learning a denoising function rather than the full time-dependent velocity field $u_{t}^{\theta}$.
+Importantly, optimizing the conditional velocity field is equivalent to optimizing the marginal velocity field [lipman2023].
+The endpoint reformulation of [dunn2024-1] reparametrizes the flow matching problem by focusing on learning a denoising function rather than the full time-dependent velocity field $u_{t}^{\theta}$.
 Specifically, the method learns a denoising function $x^{\theta}_{1|t}$ that predicts the final state $x_{1}$ from the intermediate state $x_{t}$.
@@
 where $z$ is a conditioning variable that can be either the final graph $x_{1}$ or a pair $z=(x_{0},x_{1})$ representing both the initial and final states.
-Discrete graph features are handled via CTMC as developed by Campell et al. [20] and Gat et al. [41]. In this setting, Eq. (3) is reformulated to minimize a cross-entropy loss over discrete variables:
+Discrete graph features are handled via CTMC as developed by Campell et al.
+[campbell2024] and Gat et al.
+[gat2024].
+In this setting, Eq.
+(1) is reformulated to minimize a cross-entropy loss over discrete variables:
 $\mathcal{L}_{\rm{CE}}=\mathbb{E}_{t,p_{t|1}(x_{t}|z),p_{z}}\left[-\log p^{\theta}_{1|t}(x_{1}^{i}|x_{t})\right]$ (2)
@@
 The functions $\alpha_{t}$ and $\beta_{t}$, known as the interpolant schedule, control the rate of mixing between the base and target distributions.
-A typical and effective choice of interpolant schedule for both molecule and materials generation is a linear schedule: $\alpha_{t}=(1-t)$ and $\beta_{t}=t$ [26, 18].
-Other variants of interpolants are discussed in detail in [17], and they have been validated in the context of inorganic crystals generation in our OMatG framework [18].
+A typical and effective choice of interpolant schedule for both molecule and materials generation is a linear schedule: $\alpha_{t}=(1-t)$ and $\beta_{t}=t$ [dunn2024-2, hoellmer2025].
+Other variants of interpolants are discussed in detail in [albergo2023, hoellmer2025].
 The base distribution for the atomic position flow is a standard Gaussian distribution $p_{0}(X)=\prod_{i=1}^{N}\mathcal{N}(X_{0}^{i}|\mathbf{0},\mathbb{I}_{3})$.
@@
 Interpolant for atom type, charge, and bond order.
-Atom types, charge and bond orders are modeled using CTMC-based flows [20].
+Atom types, charge and bond orders are modeled using CTMC-based discrete flow matching [campbell2024].
 For each categorical variable, a ‘mask’ token is added as an additional discrete state.
@@
 (4).
-A linear interpolant is effective for categorical variables as well [26, 18].
+A linear interpolant is effective for categorical variables as well [dunn2024-2, hoellmer2025].
 Here, $\delta(i,j)$ denotes the Kronecker delta, which is 1 if $i=j$ and 0 otherwise.
@@
 The transition probability for atom $i$ over a short time interval $\Delta t$ can be expressed in terms of individual atomic contributions:
-$p^{i}_{t+\Delta t}(j|A^{i}_{t})$ =$=\left\{\begin{array}[]{lcl}R^{i}(A^{i}_{t},j)&\mbox{for}&j\neq A_{t}^{i}\\ 1+R^{i}(A^{i}_{t},A_{t}^{i})&\mbox{for}&j=A_{t}^{i}\end{array}\right.$ (11)
+$p^{i}_{t+\Delta t}(j|A^{i}_{t})$ $=\left\{\begin{array}[]{lcl}R^{i}(A^{i}_{t},j)&\mbox{for}&j\neq A_{t}^{i}\\ 1+R^{i}(A^{i}_{t},A_{t}^{i})&\mbox{for}&j=A_{t}^{i}\end{array}\right$. (9)
 $=\delta(j,A_{t}^{i})+R^{i}(A^{i}_{t},j)\Delta t$ (10)
-where $R(A^{i}_{t},j)$ is the rate matrix specifying the transition probability from atom type $A^{i}_{t}$ to $j$ [20].
+where $R(A^{i}_{t},j)$ is the rate matrix specifying the transition probability from atom type $A^{i}_{t}$ to $j$ [campbell2024].
 To ensure proper normalization of the transition probabilities at time $t+\Delta t$, the rate matrix must satisfy $R^{i}(A^{i}_{t},A^{i}_{t})=-\sum_{j\neq A^{i}_{t}}R(A^{i}_{t},j)$ so that the total probability mass remains normalized.
 An additional stochastic term $\eta$ can be introduced into the rate matrix to form the modified rate matrix $R_{t}^{\eta}:=R_{t}^{*}+\eta R_{t}^{\mathrm{DB}}$.
-This new rate matrix generates the same marginal probability path if $R_{t}^{\mathrm{DB}}$ satisfies detailed balance over the conditional probability path $p_{t|1}$.
+This new rate matrix generates the same marginal probability path if $R_{t}^{\mathrm{DB}}$ satisfies detailed balance (‘DB’) over the conditional probability path $p_{t|1}$.
 By including the stochasticity term and choosing a linear interpolant ($\alpha_{t}=1-t$ and $\beta_{t}=t$), the rate matrix for atom $i$ becomes:
@@
 (11), the rate matrix can diverge as $t\to 1$.
-To address this issue, we employ a time-dependent loss function that avoids division-by-zero [42].
-Additionally, low-temperature sampling—which rescales the prediction logits to sharpen the output distribution—is found to be critical for improving model performance.
+To address this issue, we employ a time-dependent loss function that avoids division-by-zero [le2023].
+Additionally, low-temperature sampling—which rescales the prediction logits to sharpen the output distribution—is found to be critical for improving model performance [dunn2024-2].
 Specifically, the denoiser network’s output $p^{\theta}_{1|t}(A_{1}^{i}|A_{t})$ is transformed as follows:
@@
 During inference, we observe that turning off the stochastic term leads to better performance.
-To ensure that the generated molecules satisfy the required symmetry constraints, the joint flow matching process is constructed with an invariant base distribution and an equivariant transition probability path.
-A formal derivation of the equivariance properties of the flow matching process is provided in Supplementary Section LABEL:sec:equivariantfm.
-Property Embedding Methods
+To ensure that the generated molecules satisfy the required symmetry constraints, the joint flow matching process is constructed with an invariant base distribution and an equivariant transition probability path parameterized by the SE(3)-GVP architecture.
+A self-contained formal derivation of probability invariance and equivariant generative process is provided in respective Supplementary Section LABEL:sec:equivariantfm and Supplementary Section LABEL:sec:se3_gvp.
+Model architecture
+We utilize the FlowMol architecture, which is implemented with PyTorch and Deep Graph Library [dunn2024-2].
+Molecule updates are achieved through layers comprising Geometric Vector Perceptrons (GVP).
+Within each GVP, the molecule graph passes through a sequential steps of node feature update, node position update and edge feature update.
+Each node $i$ consists of a position $x_{i}\in\mathbb{R}^{3}$, scalar features $s_{i}\in\mathbb{R}^{d}$, and vector features $v_{i}\in\mathbb{R}^{c\times 3}$.
+The scalar feature is a concatenation of atom type and charge vectors; that is, $s_{i}:=[a_{i}:c_{i}]$ where ‘:’ defines a concatenation operation.
+Vector features are initialized at zeros, and despite not existing in the final outputs, their updates employ the vector cross-products that is not equivariant to reflections, hence SE(3) equivariant [dunn2024-1].
+Each edge feature corresponds to the bond order and the permutation invariance of the scaler bond order is ensured by taking the sum of learned bond features from $i\to j$ and $j\to i$;i.e., $\widehat{e}^{ij}=\mathrm{MLP}(e_{ij}+e_{ji})$.
+Node feature update.
+The node feature uses graph convolution to update node scalar and vector features $s_{i},v_{i}$.
+It follows two steps: the first step generates scalar $m_{i\to j}^{(s)}$ and vector messages $m_{i\to j}^{(v)}$ by a function $\psi_{M}$ which is a chain of two GVPs:
+$m_{i\to j}^{(s)},m_{i\to j}^{(v)}=\psi_{M}\left(\left[s_{i}^{(l)}:e_{ij}^{(l)}:d_{ij}^{(l)}\right],\left[v_{i}:\frac{x_{i}^{(l)}-x_{j}^{(l)}}{d_{ij}^{(l)}}\right]\right)$ (13)
+Where $d^{(l)}_{ij}$ is the distance between nodes $i$ and $j$ at update block $l$.
+In the FlowMol implementation, the distance $d_{ij}$ is replaced with a radial basis distance embedding before being fed into the GVPs or MLPs.
+Following the message generation step, an update of node scalar and vector features can be accomplished by aggregating generated messages:
+$s_{i}^{(l+1)},v_{i}^{(l+1)}=\mathrm{LN}\left([s_{i}^{(l)},v_{i}^{(l)}]+\psi_{N}\left(\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}\right)\left[m_{j\to i}^{(s)},m_{j\to i}^{(v)}\right]\right)$ (14)
+Where $LN$ stands for LayerNormalization operation, $\psi_{N}$ is a chain of three GVPs.
+Node position update.
+NPU block updates node positions by taking node-wise operations on the updated node scalar and vector features:
+$x_{i}^{(l+1)}=x_{i}^{(l)}+\psi_{P}\left(s_{i}^{(l+1)},v_{i}^{(l+1)}\right)$ (15)
+Where $\psi_{P}$ is a chain of three GVPs in which the final output has 1 vector and 0 scalar features.
+Edge feature update.
+Edge features are updated by edge-wise operations that take the updated node scalar features and node distance as the inputs:
+$e_{ij}^{(l+1)}=\mathrm{LN}\left(e_{ij}^{(l)}+\mathrm{MLP}\left(s_{i}^{(l+1)},s_{j}^{(l+1)},d_{ij}^{(l+1)}\right)\right)$ (16)
+Models are trained using the settings detailed in the ‘Hyperparameters and training settings’ section.
+Property embedding operations
 Conditional generation can be framed as sampling a property-conditional target distribution $p_{1}(x|k)$ over molecular graphs, where $k$ is a property.
-Previous approaches to conditional generation concatenate a property value to the node features of a molecular graph, training the model jointly on the property and the graph [7, 8, 21].
+Previous approaches to conditional generation concatenate a property value to the node features of a molecular graph, training the model jointly on the property and the graph [hoogeboom2022, xu2023, song2023].
 As a result, the learned denoising process is conditioned on the property.
@@
 The dimensionality of the property embedding matches that of the node scalar features, allowing seamless integration without requiring additional dimensional matching.
-To encode interactions between the property embedding and node features, we propose five distinct operations.
+To encode interactions between the property embedding and node scalar features in molecualr graphs, we propose five distinct operations, including ‘Concatenation’, ‘Sum’, ‘Multiply’, ‘Concatenate + Sum’, and ‘Concatenate + Multiply’.
 Depending on the operation, if the combined feature vector $[A,C]$ maintains its original dimensionality after interacting with the property embedding, no further processing is applied; otherwise, an MLP is used to project it back to the original dimensionality.
-Since the property embedding interacts exclusively with node scalar features, leaving position vectors unchanged, the resulting property-conditioned graph retains SE(3) equivariance.
+Since the property embedding interacts exclusively with node scalar features, leaving position vectors unchanged, the resulting property-conditioned graph retains the SE(3) equivariance.
 Moreover, the loss function preserves its original form, and the trained SE(3) GNN defines a valid conditional generative process.
-In the ‘Sum’ operation, the property embedding is simply added to the node scalar features.
+Specifically, in the ‘Sum’ operation, the property embedding is simply added to the node scalar features.
 The ‘Multiply’ operation performs an element-wise Hadamard product $\odot$’.
 To ensure the multiplicative factors remain within a bounded range, the property embedding is passed through a sigmoid function and then shifted by 0.5, resulting in multipliers in the interval $[0.5,1.5]$.
-Both ‘Sum and Multiply’ preserve the original feature dimensionality.
+Both ‘Sum’ and ‘Multiply’ preserve the original feature dimensionality.
 In contrast, operations involving ‘Concatenate’ increase the dimensionality of the node scalar features, and an MLP transformation $\varphi_{\theta}$ is required to project the combined features back to the original dimension.
@@
 In both cases, the final output has the desired dimensionality, either by design or via the subsequent MLP transformation.
-Gaussian Expansion for Augmented Property Embedding
-We incorporate a non-trainable Gaussian expansion layer to enrich the property representations prior to being mapped to a property embedding via an MLP $\phi_{\mathrm{GE}}$ [29].
+Gaussian expansion
+We incorporate a fixed, non-trainable Gaussian expansion layer to enrich the property representations prior to being mapped to a property embedding via an MLP $\phi_{\mathrm{GE}}$ [gebauer2022].
 The Gaussian expansion maps a scalar property value into a fixed-length vector using a set of Gaussian basis functions.
@@
 The top panel of Figure 1d shows the five Gaussian basis functions, with their centers marked by vertical dashed gray lines.
-In the bottom panel, we pick three example property values, namely 0.1, 0.3, and 0.5 Hartree, and show the expanded feature vectors as bar plots indicating the activation of each Gaussian basis function.
+In the bottom panel, we pick two example property values, namely 0.1, and 0.5 Hartree, and show the expanded feature vectors as bar plots indicating the activation of each Gaussian basis function.
 For instance, if the gap is 0.1 Hartree, the first two Gaussian functions (centered closest to 0.1) have higher responses than the remaining functions, whose centers are farther away.
 Similar activation patterns are observed for the other two property values.
-Datasets and Baselines
-Datasets.
-We used RDKit’s sanitization function [33] to filter chemically valid molecules on the QM9 SDF file with bond and charge fix, resulting in a curated set of 133k molecules.
-After random shuffling, we split the dataset into 100k training, 20k validation, and 13k test samples.
+Inference
+ID tasks.
+To generate paired samples of property values and numbers of atoms, we employed a two-stage sampling procedure.
+First, atom counts were sampled according to their empirical distribution in the training dataset.
+Specifically, the frequency of each unique atom count was computed and used to construct a categorical probability distribution, from which atom counts were drawn.
+Conditional on a given atom count, the associated property values were sampled using a histogram-based approximation of their empirical distribution.
+For each unique atom count, the observed property values were discretized into a fixed number of bins like 1000.
+The counts within each bin were normalized to form a categorical distribution.
+During sampling, a bin index was first drawn from this categorical distribution, after which a continuous property value was assigned by uniformly sampling within the range of the selected bin.
+This approach yields paired samples of atom counts and property values that respect both the marginal distribution of atom counts and the conditional distribution of properties given the atom count.
+To make a fair comparison across different property embedding methods, the same sampled target values and atom counts were used.
+OOD tasks.
+The property values in the OOD tasks are chosen as the 99th quantile of the QM9 training data and the numbers of atoms are selected for the property ranges bounded by the 97th quantile and the maximum (100th quatile) property value.
+During inference, PropMolFlow models use Euler’s method with 100 evenly spaced time steps to integrate the learned velocity field.
+Datasets, baselines and evaluation metrics
+Datasets
+We trained PropMolFlow on the QM9 data with explicit hydrogen atoms [ramakrishnan2014quantum, wu2018moleculenet].
+We did not evaluate PropMolFlow on the GEOM-Drugs dataset, which is commonly used for unconditional generation and includes larger molecules, because it only provides quantum-mechanical energy values and lacks other relevant properties [axelrod2022].
+The original QM9 data contains 133,885 molecules, and each molecule contains up to 9 heavy atoms and consists of 3–29 atoms (average: 18 atoms), composed of up to 5 elements: C, N, O, F, and H.
+Molecules are fully optimized and molecular properties are obtained with DFT at the level of B3LYP/6-31G(2df,p) [becke1993, lyp1988].
 For property-guided generation, we considered six molecule properties: polarizability ($\alpha$), HOMO-LUMO gap, HOMO energy, LUMO energy, dipole moment ($\mu$), and heat capacity ($C_{v}$).
 Precise definitions of each property can be found in Supplementary Section LABEL:sec:more_details.
-We did not evaluate PropMolFlow on the GEOM-Drugs dataset, which is commonly used for unconditional generation and includes larger molecules, because it only provides quantum-mechanical energy values and lacks other relevant properties [43].
+Upon inspection, we discovered numerous inconsistencies: invalid bond orders and non-zero net charges in roughly 30,000 molecules, despite QM9’s requirement for charge-neutral, closed-shell valency [ramakrishnan2014quantum].
+We corrected these discrepancies by reassigning bond orders or charges to enforce valency-charge consistency for the vast majority of entries and used the corrected data to train our PropMolFlow models.
+After the correction, we improved the structural validity of generated molecules substantially (Supplementary Table LABEL:tab:data_and_metric_compare).
+Previous works that used bond orders overlooked this issue [vignac2023, dunn2024-2]; instead, they modified evaluation metrics for models trained on the problematic SDF data against the one used by Hoogeboom et al.
+[hoogeboom2022] to tolerate chemically inconsistent bond–charge combinations, hence inflating molecule stability estimates contrary to QM9’s intended constraints.
+A detailed comparison between models using previous problematic data and the corrected data, together with the comparison between previous stability metrics and the revised ones, can be found in Supplementary Table LABEL:tab:data_and_metric_compare.
+A detailed description of our procedure and the alignment with the original QM9 XYZ data is provided in Supplemental Section LABEL:sec:qm9_data_fix.
+The revised QM9 SDF data is provided on Zenodo [zenodo_propmolflow].
+We used RDKit’s sanitization function [rdkit] to filter chemically valid molecules on the QM9 SDF file with bond and charge fix, resulting in a curated set of 133k molecules.
+After random shuffling, we split the dataset into 100k training, 20k validation, and 13k test samples.
 PropMolFlow models and GVP property predictors are trained on disjoint 50k and 50k datasets from the 100k training data.
-Baselines.
-We compare PropMolFlow against several recent property-guided molecular generation methods based on diffusion models and flow matching, including EEGSDE, GCDM, GeoLDM, JODO, and EquiFM [27, 25, 8, 44, 21].
+Baselines
+We compare PropMolFlow against several recent property-guided molecular generation methods based on diffusion models and flow matching, including EEGSDE, GCDM, GeoLDM, JODO, and EquiFM [bao2023, morehead2024, xu2023, huang2024, song2023].
+EEGSDE (Equivariant Energy Guided Stochastic Differential Equations) use an energy function to guide the diffusion for property-guided generation [bao2023].
+EquiFM is the first equivariant flow matching model for 3D molecule generation [song2023].
+GeoLDM represents the first latent diffusion model for molecular geometry generation [xu2023].
+GCDM constructs an SE(3) equivariant geometry-complete diffusion model using geometry-complete perceptrons for molecule generation [morehead2024, moret2023].
+JODO uses a diffusion graph transformer for joint generation of 2D molecular graphs and 3D geometries.
+EEGSDE, EquiFM, GeoLDM, and GCDM do not include bond orders in their molecular graphs, while JODO incorporates both bond existence and bond orders.
+All five baseline models omit atomic charges in their conditional generation, in contrast to PropMolFlow, which can generate molecules with non-zero atomic charges.
 In this work, we focus exclusively on conditional generation results; unconditional generation performance is reported in earlier studies.
 Language-model-based approaches are excluded from this comparison, as state-of-the-art models in that category rely on fundamentally different molecule representations and training data regimes.
-Our comparison includes both structural-validity and property-specific evaluation metrics.
-For EEGSDE, we benchmark PropMolFlow against the canonical conditional model with a scaling factor of 1 ($s=1$[27].
+For EEGSDE, we benchmark PropMolFlow against the canonical conditional model with a scaling factor of 1 (s=1s=1[bao2023].
 Although EEGSDE provides DFT-based validation for five out of six properties, it does so on only 100 generated molecules, which limits the statistical significance of those results.
-EEGSDE reported the atomic and molecular stability results for its conditional generation, JODO reported molecular stability, and GCDM reported PoseBusters validity for the conditional generation of GCDM and GeoLDM models—no other baseline conditional models provided these metrics.
-Because EEGSDE and GCDM released their model checkpoints, we sampled from those and computed all structural validity metrics on the resulting structures; for GeoLDM, we used the checkpoints bundled with GCDM.
-EquiFM checkpoints were unavailable, so we trained its conditional models ourselves following information in [21] (training details in Supplementary Section LABEL:sec:more_details).
+Because EEGSDE, JODO and GCDM released their model checkpoints, we sampled molecules from those and computed all structural validity metrics on the resulting structures; for GeoLDM, we used the checkpoints bundled with GCDM.
+EquiFM checkpoints were unavailable, so we trained its conditional models by ourselves following information in [song2023] (training details in Supplementary Section LABEL:sec:more_details).
+Three intuitive baselines are included for comparison in the ID task.
+The “Random (Upper-bound)” intuitive baseline fully shuffles property values, thereby removing any correlation between molecular structures and properties, and serves as an upper bound on error.
+The “# Atoms” baseline uses only the number of atoms as a predictor for molecular properties, capturing coarse size-dependent trends.
+It is implemented as a simple neural network with one hidden layer [hoogeboom2022].
+The “QM9 (Lower-bound)” baseline corresponds to a pretrained EGNN regressor provided by Hoogeboom et al.
+[hoogeboom2022], trained directly on QM9 and serving as a lower bound on achievable error.
+Improvement over the “Random” baseline indicates that the conditional generation effectively integrates property information.
+Surpassing the “# Atoms” baseline suggests that the generative model captures structural features beyond simple atom count when generating new molecules.
+Evaluation metrics
+Structural validity metrics.
+Our comparison includes both structural-validity, inference-efficiency and property-specific evaluation metrics.
+Structural validity metrics include atomic stability, molecule stability, RDKit validity, ‘uniqueness and validity’, PoseBusters validity, and closed-shell ratio.
+Atom stability is given by the proportion of atoms with correct valency.
+To account for atomic charges, an atom is considered to have the correct valency if its formal charge balances its explicit valency, as defined in RDKit [rdkit].
+For instance, a nitrogen atom should carry a +1 formal charge if it has a valency of 4.
+Molecule stability defines the proportion of molecules in which all atoms are stable and the molecule is charge neutral (zero net formal charges).
+RDKit validity refers to the ratios of molecules that pass RDKit’s sanitization check.
+‘Uniqueness and validity’ is defined as the proportion of molecules that are RDKit valid and unique in their SMILES representation.
+PoseBusters validity corresponds to the fraction of molecules that pass many de novo chemical and structural validation tests, such as all atom connectivity, valid bond lengths and angles, and absence of internal clashes [buttenschoen2024].
+We used the first 10 columns of the PoseBusters outputs in CSV formats as the criteria for results shown in Fig.
+2 and Supplementary Table LABEL:tab:standard_metrics.
+A molecule is closed-shell valid if it has an even number of valence electrons.
+Since all molecules in QM9 are closed-shell, any generated open-shell species are considered invalid.
 Except for the closed-shell ratio, all structural validity metrics depend on bond-order information, which baseline models (other than JODO) do not output.
-We therefore assigned bond orders using the distance-based cutoffs introduced by Hoogeboom [7].
-Although recent work [23] suggests that the bond orders can probably be optimized through the OpenBabel program [45].
-We did not apply such refinements to the baseline samples.
-For property metrics, PropMolFlow is directly compared to prior methods on the ID tasks.
-OOD results are not directly comparable due to a lack of baseline generative models and property prediction models testing this regime, but we validate our results with DFT calculations to assess physical fidelity.
+We therefore assigned bond orders using the distance-based, optimized cutoffs introduced by Hoogeboom [hoogeboom2022].
+Although recent work [vignac2023] suggests that the bond orders can probably be optimized through the OpenBabel program [O'Boyle2011], we did not apply such refinements to the baseline samples.
+To assess the chemical similarity between generated molecules and a reference training set, we employed a fingerprint-based approach using RDKit.
+For each molecule, a Morgan fingerprint (circular fingerprint) was computed with a radius of 2 and a 2048-bit representation.
+Pairwise similarity between each generated molecule and all training molecules was computed using the Tanimoto similarity metric on their respective Morgan fingerprints.
+For each generated molecule, the maximum similarity score and the corresponding most similar training molecule were recorded.
+Lastly, inference efficiency is given by the wall-clock time required to generate 10,000 molecules.
+Property metrics.
+PropMolFlow is directly compared to prior methods on the ID tasks.
+In this task, a conditional generative model was first used to generate molecules conditioned jointly on property values and atom counts sampled from the same distribution as the training data (for example, QM9).
+These sampled property values and atom counts served as the input conditions.
+A separate property predictor/regressor was then applied to estimate the properties of the generated molecules.
+In PropMolFlow, we used an SE(3) GVP network for molecule generation and, correspondingly, trained a separate GVP regressor for each of the six molecular properties.
+10000 molecules are generated to evaluated the performance of conditional generation.
+OOD results are not directly comparable to baseline models, but we validate our results with DFT calculations to assess physical fidelity.
+RMSD and property sensitivity.
+Given the same chemical compositions and element-type ordering, the RMSD between any two molecules (M0 and M1) is the minimum distance between these two structures considering the translational and rotational symmetry operations, which can be expressed as:
+$\mathrm{RMSD}(\rm{M}_{0},\rm{M}_{1})=\min_{P\in SE(3)}\{d\left(M_{0},P(M_{1})\right)\}$ (18)
+Where SE(3) represents the group of operations that respect the translational and rotational symmetry.
+To estimate the property sensitivity to structural relaxation, we selected the structures whose RMSDs are larger than 0.03 Å—RMSDs lower than this threshold are considered very close to the relaxed structures, and in Eq.
+(19), we define the normalized property sensitivity $\chi_{\rm{DFT}}$’ to allow comparisons between different properties.
+$\chi_{\rm{DFT}}=\frac{|{\delta q_{\rm{DFT}}|}}{(q_{0.99}-q_{0.01})\cdot\rm{RMSD}}$ (19)
+where $\delta q_{\rm{DFT}}$ is the change of DFT-calculated property upon relaxation, and the normalizer ($q_{0.99}-q_{0.01}$) is the difference between the 99th quantile and the 1th quantile of the property distribution of QM9 training data.
+This range is chosen over the min-max range because it is robust to outliers.
 Computational settings
-Hyperparameters, Training and Inference Details
-PropMolFlow uses the model architecture detailed in Supplemental Section LABEL:sec:appendix_model.
+GVP regressor details
+To be self-consistent, we trained property regressors using Graph Neural Networks based on GVPs.
+GVP property predictors are trained on a disjoint dataset versus that used for training PropMolFlow models, hence avoiding data leakage between the generative models and property prediction models.
+We appended an MLP layer that takes the final node scalar features as input to predict the target property.
+The parameters of GVP regressors are optimized by minimizing a mean squared error loss function.
+Separate models were trained for each of the six molecular properties.
+DFT
+All DFT calculations were performed using the Gaussian 16 package (Gaussian 16, Revision C.01) [g16].
+The B3LYP hybrid functional [becke1993, lyp1988] together with the 6-31G(2df,p) basis set was employed for single-point calculations, and geometric optimizations.
+Single-point calculations were used unless otherwise specified.
+Structural relaxations were performed for the property $C_{v}$ in most cases, except in the interpolation study.
+Relaxations were performed to evaluate changes in both properties and geometries, as shown in Fig.
+3.
+DFT calculations that failed to converge were discarded.
+We released the DFT-evaluated molecules in both the ID and OOD tasks with and without structural optimization at the Zenodo repository [zenodo_propmolflow].
+Molecule filtering
+We selected 1000 molecules out of 10000 samples structures in the ID task or use the directly generated 1000 structures in the OOD task for DFT evaluations.
+multi-step filtering is essential to ensure that the structures are chemically valid and suitable for downstream applications.
+We filtered the 1000 molecules by five criteria: Molecule stability, RDKit validity, PoseBusters validity, closed-shell validity, and multi-fragment check.
+The multi-fragment check filters out molecules that with more than one disconnected fragment.
+For molecules passing all filters, property values are evaluated using both a GVP property predictor and DFT calculations.
+We also filtered out molecules whose DFT calculations did not converge, resulting in no less than 90% of and 78–94% of the molecules remaining for the subsequent analyses in the respective ID and OOD tasks (Supplementary Table LABEL:tab:no_sp_dft).
+Hyperparameters and training details
 PropMolFlow models on the QM9 dataset were trained with 8 molecule update blocks.
@@
 QM9 PropMolFlow models were trained with 2000 epochs.
-Training and inference for PropMolFlow models and all baseline models used a single NVIDIA A100-SXM4 graphic card with 80GB of memory, with a batch size of 128.
-PropMolFlow models can be trained in about 4–5 days, and EquiFM models were trained with 2500 epochs, which took around 3.5 days.
-During inference, PropMolFlow models use Euler’s method with 100 evenly spaced time steps to integrate the learned velocity field.
-The number of atoms for each generated molecule is first sampled from the QM9 atom count distribution, and the property values are then sampled according to the number of atoms to ensure consistency with the dataset’s statistics, as used in previous property-conditioning works [7, 8, 25].
-GVP regressor details.
-To be self-consistent, we trained property regressors using Graph Neural Networks based on GVPs.
-GVP property predictors are trained on a disjoint dataset versus that used for training PropMolFlow models, hence avoiding data leakage between the generative models and property prediction models.
-We appended an MLP layer that takes the final node scalar features as input to predict the target property.
-The parameters of GVP regressors are optimized by minimizing a mean squared error loss function.
-Separate models were trained for each of the six tested properties.
-Acknowledgment
+For each property embedding method, 6 model checkpoints with lowest validation losses were saved and the top-3 model checkpoints were deposited to the Zenodo repository [zenodo_propmolflow].
+Inference for PropMolFlow models and all baseline models used a single NVIDIA A100-SXM4 graphic card with 80GB of memory, with a batch size of 128.
+Training of PropMolFlow models use both Nvidia A100-SXM4 and 2080ti graphic cards.
+PropMolFlow models can be trained in 2–3 days with A100 and 4–5 days with 2080Ti cards, and EquiFM models were trained with 2500 epochs on A100, which took around 3.5 days.
+Data availability
+All data can be found in the Zenodo repository [zenodo_propmolflow], including revised QM9 SDF data, SDF files for generated raw molecules, DFT-calculated structures and their molecular properties saved in the extxyz format, full property MAE and structural validity results in CSV format for all saved PropMolFlow checkpoint models, sampled structures and their full PoseBusters results for all baseline models, and EGNN property-predictor pretrained models and notebook examples to use the property predictors.
+Code availability
+Our PropMolFlow implementation is available at https://github.com/Liu-Group-UF/PropMolFlow and on Zenodo [propmolflow_code].
+Acknowledgments
 The authors acknowledge funding from NSF Grant OAC-2311632 and from the AI and Complex Computational Research Award of the University of Florida.
@@
 We would like to thank Alex Morehead (University of Missouri) for providing the checkpoint models for the conditional generation of GeoLDM.
-Data and code availability
-Code and data of this work can be found on github: https://github.com/Liu-Group-UF/PropMolFlow.